### v6

# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

Abstract
This paper establishes the first- and second-order asymptotics of the maximal throughput under a covertness criterion defined by a Kullback-Leibler (KL) divergence bound $\delta$. The achievability employs random coding with three key innovations: (1) dynamically scaled power constraints dependent on blocklength $n$, (2) codeword generation via truncated Gaussian distributions optimized for near-isotropy, and (3) codebook expurgation to ensure reliability. The analysis reveals two critical results. First, compared with the previous results, we present analytic formulas of the first-and second order asymptotics as functions of blocklength $n$, decoding error probability $\epsilon$, the adversary’s noise variance $\sigma^{2}$ and the upper bound $\delta$ on KL divergence. Second, the required secret key size exhibits a universal scaling law of $O(n^{\frac{1}{2}})$ even when the legitimate receiver’s channel quality marginally exceeds the adversary’s — a sharp contrast to discrete memoryless channels where such asymmetry reduces key requirements. Theoretical contributions are further reinforced through an information-geometric perspective: truncated Gaussians minimize detectability by confining power within $O(n^{\frac{1}{2}})$ while maximizing throughput through near-capacity-achieving properties. The square-root power law is further justified via asymptotic analyses of several divergences of Gaussian family with different variances. These results advance covert communication design by providing explicit scaling laws, optimal codebook architectures, and a systematic methodology for balancing reliability, covertness, and resource constraints.
I Introduction
Security is a very important aspect of wireless communications. Covert communication or communication with low probability of detection (LPD), which requires that the adversary not be able to detect whether the legitimate parties are communicating, has been studied in many recent works [1] - [8]. Typical scenarios include underwater acoustic communication [9] and dynamic spectrum access in wireless channels, where secondary users attempt to communicate without being detected by primary users or users wish to communicate while avoiding attention from regulatory entities [10]. The information theory for covert communications was first characterized for AWGN channels in [11] and for discrete memoryless channels (DMCs) in [10][12], and later in [13] and [14] [15] [16] for BSC and MIMO AWGN channels, respectively. It has been shown that the first order asymptotics of the maximal throughput of covert communication follows Square Root Law, which states that no more than $O(\sqrt{n})$ bits can be communicated over $n$ channel uses of AWGN channels or discrete memoryless channels under the condition that the adversary cannot effectively detect the presence of the communication. The information theoretical limits of covert communication has been extended in [17] for a non-coherent fast Rayleigh-fading wireless channel and in [18] for a continuous-time Poisson channel, and in [19] over quantum channels, respectively. In these works, the focus is on the first order asymptotics and the covert constraint is given as an upper bound of KL divergence between the output distributions with and without the presence of communication.
Considering the latency and practical length of the codewords, the finite blocklength performance [20] is important for practical performance evaluation. The characterization on the second order asymptotics of the maximal throughput over covert channels is necessary in a finite blocklength analysis, which represents the back-off quantity with given finite blocklength $n$, probability $\epsilon$ of decoding error and some covert metric $\delta$ from the first order asymptotics. Tahmasbi and Bloch characterize the second order asymptotics of covert communication over discrete memoryless channels in [21][22] from the perspective of channel resolvability. A closed form of the second order asymptotics is given when the covert metric is KL divergence, and related bounds are provided when the covert metrics are total variation distance (TVD) and missed detection probability, respectively. In [7] and [23], one-shot achievability bound of Gaussian random coding over AWGN channels under maximal power constraint are presented together with the TVD between the distributions of channel outputs at the adversary with and without communications of the legitimate transmitter. The converse bound under a maximal power constraint that is varying with blocklength $n$ is also provided. Nevertheless, the characterization of the first and second order asymptotics as a function of the covert constraint $\delta$ over AWGN channels is still missing.
This paper investigates the first- and second-order asymptotics of covert communication over AWGN channels, characterized as functions of blocklength $n$ decoding error probability $\epsilon$, the noise level at the adversary $\sigma$ and covertness parameter $\delta$, herein $\delta$ constrains the Kullback-Leibler (KL) divergence between the adversary’s observation distributions with and without communication. For achievability analysis, we extend our previous framework in [23] by maintaining the truncated Gaussian generating distribution with hyper-parameter $\mu$, ̵while replacing the sequential encoding scheme with an enhanced approach combining i.i.d. sampling and expurgation operations from [22]. Our reliability analysis introduces a novel decoding threshold selection mechanism that effectively bounds error probabilities for codewords with varying radii – ̵a distinctive feature not addressed in prior works [20, 24]. The channel resolvability analysis advances in two key aspects: First, ̵we circumvent the quasi-metric limitations of KL divergence in Euclidean space by directly analyzing the divergence between the Gaussian mixture induced by codewords and the noise distribution, avoiding reliance on modified triangle inequalities applicable to DMCs. Second, ̵we establish an explicit connection between adversary noise levels and required key bits, revealing a fundamental dichotomy between AWGN and DMC scenarios. ̵Notably, the key size remains $O(\sqrt{n})$ even when adversary noise marginally exceeds main channel noise – a critical divergence from DMC behavior that we rigorously explain through channel characteristic analysis. The optimality of truncated Gaussian distributions is justified through local information geometry principles [26, 27], where covertness realization corresponds to positioning the signal distribution within a quasi-$\varepsilon$ neighborhood of noise distribution in statistical manifold space. ̵This geometric interpretation bridges coding theory with machine learning applications [28, 29]. The square root law governing power scaling $\Psi(n)=O(\frac{1}{\sqrt{n}})$ emerges naturally from asymptotic analyses of Gaussian divergence measures and noise distribution approximation accuracy. Note that our work only considers AWGN channel model and its performance limit under finite blocklegnth regime. We do not take into account of modulation, the signal duration, or spectral width. Under such circumstance, the transmission involves specific modulation, and has been investigated in [30][31].
Our principal contributions include:
  • Complete asymptotic characterization: First- and second-order expansions for Gaussian random coding in AWGN covert communication, explicitly parameterized by $(n,\epsilon,\sigma,\delta,\mu)$. ̵The $\mu$-dependence underscores finite-blocklength power constraint necessities.
  • Key size quantification: Determination of minimal key requirements where first-order asymptotics depend on adversary noise levels. ̵The persistent $O(\sqrt{n})$ scaling despite favorable noise conditions reveals intrinsic AWGN-DMC differences.
  • Information geometric justification: Unification of truncated Gaussian optimality with local neighborhood structures in statistical manifolds.
  • Divergence-driven power law: Rigorous explanation of square root law through asymptotic Gaussian divergence behavior and noise approximation theory.
The rest of this paper is organized as follows. In Section II, we present some preliminaries on notations, the channel model, the outline of the proof of the achievability and the converse, the definitions of codes and some distributions. Section III presents the main results, which include the first and second order asymptotics of the achievability bound and the investigation of the converse bound. The framework of local information geometry to explain the optimality of truncated Gaussian distribution, the investigation on the connection between Square Root Law, the choice of the power level and the asymptotic behavior of several divergences of Gaussian distributions are provided in Section IV. Section V concludes the paper.
II Preliminaries
II-A Notations
Throughout the paper, $\log$ denotes logarithmic function with base $2$. We use upper case letters, such as $X$, for random variables and lower case letters, such as $x$ for their realizations. Moreover, upper case letters with boldface fonts are used for random vectors, e.g., $\bm{Z}$, and their realizations are denoted with lower case letters of boldface font, e.g., $\bm{z}$. The norm $\|\cdot\|$ will be used for $l^{2}$ norm $\|\cdot\|_{2}$. We use * for convolution operation and $\langle M\rangle$ for the set $\{1,\cdots,M\}$. If $a$ is a real number, then $|a|$ denotes its absolute value. If $\mathcal{X}$ is a set, then $|\mathcal{X}|$ denotes its cardinality. The function $\Phi$ denotes the cumulative distribution function of standard normal random variable. For two distributions $P$ and $Q$ in the same probability space, some related divergences in this paper are defined as follows.
The Kullback-Leibler divergence (KL divergence)
$\mathbb{D}(P\|Q)\triangleq\int\log\frac{dP}{dQ}dP.$ (1)
The total variational distance (TVD)
$\mathbb{V}(P,Q)=\frac{1}{2}\int|\frac{dP}{dQ}-1|dQ.$ (2)
The square of the Hellinger distance
$H^{2}(P,Q)=\frac{1}{2}\int\big(\sqrt{dP}-\sqrt{dQ}\big)^{2}.$ (3)
$\chi^{2}$-divergence:
$\mathbb{\chi}^{2}(P\|Q)\triangleq\int\big(\frac{dP}{dQ}-1\big)^{2}dQ.$ (4)
II-B The Channel Model
The AWGN covert channel model considered in this work is illustrated in Fig.1. A transmitter Alice communicates with a legitimate receiver Bob through a main channel $W_{1}$, and the signals are also observed by an adversary Willie through another channel $W_{2}$, which are modeled as
$W_{1}(y|x)=\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{(y-x)^{2}}{2}\right),$ (5)
$W_{2}(z|x)=\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{(y-x)^{2}}{2\sigma^{2}}\right).$ (6)
The messages are indexed by $w\in\langle M\rangle$. A codeword over the AWGN channel is represented by a vector in $\mathbb{R}^{n}$: $\bm{x}=(x_{1},x_{2},\cdots,x_{n})$. With channel transition models $W_{1}(y|x)$ and $W_{2}(z|x)$, we have $W^{n}_{1}(y^{n}|x^{n})=\prod_{i=1}^{n}W_{1}(y_{i}|x_{i})$ and $W^{n}_{2}(z^{n}|x^{n})=\prod_{i=1}^{n}W_{2}(z_{i}|x_{i})$.
II-C Definitions for Codes over AWGN Covert Channels
First, we provide formal definitions for codebooks over AWGN channels as follows.
Definition 1.
A $(M,n,\epsilon,\Psi)_{m}$ code $\mathcal{C}=\{\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{M}\}$ for a AWGN channel (5), consists of a message set $W\in\langle M\rangle$, an encoder at the transmitter Alice $f_{n}:\mathcal{W}\rightarrow\mathbb{R}^{n},w\mapsto\bm{x}$, and a decoder at the legitimate receiver Bob $g_{n}:\mathbb{R}^{n}\rightarrow\langle M\rangle,\bm{y}\rightarrow\hat{w}$ with the condition that the encoder and decoder pair satisfies
  1. 1.
    $P_{err}\triangleq\max_{w\in\mathcal{W}}\mathbb{P}\left[\hat{w}\neq w|W=w\right]\leq\epsilon.$ (7)
  2. 2.
    maximal power constraint: each codeword $\bm{x}_{i}$ satisfies
    $\|\bm{x}_{i}\|^{2}\leq n\Psi,$ (8)
If the maximal power constraint is substituted by average power constraint: the codebook satisfies
$\frac{1}{M}\sum_{i=1}^{M}\|\bm{x}_{i}\|^{2}\leq n\Psi,$ (9)
then it is a $(M,n,\epsilon,\Psi)_{a}$ code.
Fig. 1: The model of covert communication over AWGN channels in Section II
For AWGN covert channels, the following formal definition characterizes the codes applicable for them.
Definition 2.
An $(M,K,n,\epsilon,\sigma,\delta)$ code for AWGN covert channels (5) and (6) consists of an encoder/decoder pair
$f:\langle K\rangle\times\langle M\rangle\rightarrow\mathbb{R}^{n}\,\ \ \,\text{and}\,\ \ \ \ \,\phi:\mathbb{R}^{n}\times\langle K\rangle\rightarrow\langle M\rangle,$
with
$\bm{x}_{sw}=f(s,w)\,\ \ \,\text{and}\,\ \ \ \ \,\hat{w}=\phi(\bm{y},s)$ (10)
The codewords $\bm{x}_{sw},s\in\langle K\rangle,w\in\langle M\rangle$ compose a codebook satisfies the following properties.
  1. 1.
    Each message $w\in\langle M\rangle$ should be transmitted reliably:
    $P_{err}\triangleq\underset{s\in\langle K\rangle,w\in\langle M\rangle}{\max}\mathbb{P}(\phi(\bm{Y},s)\neq W|S=s,W=w)\leq\epsilon,$ (11)
  2. 2.
    The adversary should have low probability on successful detection of the communication, which leads to the KL divergence requirement on the distribution $\hat{Q}^{(n)}_{\bm{Z}}$ with respect to $\hat{Q}^{(n)}_{0}$:
    $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\leq\delta,$ (12)
    where $Q^{(n)}_{\bm{Z},\bm{C}}$ is a mixture Gaussian distribution induced by the codebook at the adversary:
    $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})=\sum_{s=1}^{K}\sum_{w=1}^{M}\frac{1}{MK}W^{n}_{2}(\bm{z}|\bm{x}_{sw}).$ (13)
    and $Q^{(n)}_{\bm{Z},0}$ is the distribution of channel noise at the adversary with blocklength $n$: $Q^{(n)}_{\bm{Z},0}(\bm{z})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}_{n})$.
Note that from Definition 1, there is always power constraint for codebooks over AWGN channels. For codes for the AWGN covert channel, there are two additional ingredients. (1), The noise level $\sigma$ of the channel $W_{2}$ plays an important role and will affect the power level of legitimate codes because the covert constraint (12) depends on the parameter $\sigma$; (2), To accomplish the target of covertness, some shared secret key $S$ may be necessary. To take the power level of codes into consideration, we introduce the following extension for AWGN covert codes by combining Definition 1 and Definition 2.
Definition 3.
  • An $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$ code is a $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$ code if for each $s\in\langle K\rangle$, the sub-code $\{\bm{x}_{sw},w\in\langle M\rangle\}$ satisfies maximal power constraint:
    $\|\bm{x}_{sw}\|^{2}\leq n\Psi\,\ \,\text{holds for each}\,\ \,w\in\langle M\rangle$ (14)
  • An $(M,K,n,\epsilon,\Psi,\sigma)_{a}$ code is a $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{a}$ code if for each $s\in\langle K\rangle$, the sub-code $\{\bm{x}_{sw},w\in\langle M\rangle\}$ satisfies average power constraint:
    $\frac{1}{M}\sum_{w=1}^{M}\|\bm{x}_{sw}\|^{2}\leq n\Psi.$ (15)
The maximal sizes of $(M,n,\epsilon,\Psi)_{m}$, $(M,n,\epsilon,\Psi)_{a}$, $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$, $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{a}$, $(M,K,n,\epsilon,\sigma,\delta)$ codes are denoted by $M^{*}_{m}(n,\epsilon,\Psi)$, $M^{*}_{a}(n,\epsilon,\Psi)$, $M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)$, $M^{*}_{a}(n,\epsilon,\Psi,\sigma,\delta)$ and $M^{*}(n,\epsilon,\sigma,\delta)$ respectively. For AWGN covert channels, we are interested in the quantity $M^{*}(n,\epsilon,\sigma,\delta)$ as a function of $n,\epsilon,\sigma,\delta$ since the covert constraint (12) always implies power constraint on codewords. In addition, for any $(M,K,n,\epsilon,\sigma,\delta)$ code, we want to characterize the corresponding number of key bits, which is denoted as $K(n,\epsilon,\sigma,\delta)$.
Remark 1.
In this work, we follow our previous works [7][23] and [22], and choose the maximal probability of error as the reliability metric.
Remark 2.
The detection performance of the adversary is usually measured by $\mathbb{V}(Q^{(n)}_{\bm{Z},\bm{C}},Q^{(n)}_{\bm{Z},0})$ [34]. When it is close to 0, it is generally believed that no detector at Willie can effectively discriminate the induced output distribution and the distribution of noise, thus cannot distinguish whether Alice is communicating with Bob. Nevertheless, it is more convenient to use KL divergence to measure the discrimination between $Q^{(n)}_{\bm{Z},\bm{C}}$ and $Q^{(n)}_{\bm{Z},0}$ in the asymptotic setting as a metric of covertness. In this work, we follow the conventions in the works [11][12] to use (12) as covert constraint.
Remark 3.
As the distribution $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})$ depends on the codewords $\bm{x}_{sw}$, which are randomly generated, the quantity $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})$ itself is a random variable.
For the above quantities, we have the following lemma, the proof of which can be found in Appendix A.
Lemma 1.
$M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)\leq M^{*}_{m}(n,\epsilon,\Psi)\leq M^{*}_{a}(n,\epsilon,\Psi),$ (16)
$M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)\leq M^{*}(n,\epsilon,\sigma,\delta),$ (17)
$M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)\leq M^{*}_{a}(n,\epsilon,\Psi,\sigma,\delta)\leq M^{*}(n,\epsilon,\sigma,\delta).$ (18)
The aim of this work is to characterize the quantity $M^{*}(n,\epsilon,\sigma,\delta)$ through the following two parts.
  1. 1.
    Achievability: A lower bound of $M^{*}(n,\epsilon,\sigma,\delta)$ is obtained by constructing a code satisfying (11) and (12) by some coding scheme, usually through random coding argument. Meanwhile, the corresponding number of key bits is denoted as $K(n,\epsilon,\sigma,\delta)$.
  2. 2.
    Converse: an upper bound of $M^{*}(n,\epsilon,\sigma,\delta)$ is provided such that any code satisfying (11) and (12) should have size larger than it.
II-D Main Results and Outline of The Proof
In this section, we provide the main results and a road map on the proof of the achievability and converse bounds.
Theorem 1.
For AWGN covert channels with channel models in (5) and (6), we have
$\log$ $M^{*}(n,\epsilon,\sigma,\delta)\geq\mu\sigma^{2}(\delta n\log e)^{1/2}$
$+$ $\sigma(2\mu)^{1/2}(n\delta)^{1/4}(\log e)^{3/4}\Phi^{-1}(\epsilon)+o(n^{\frac{1}{4}}).$ (19)
and
$\log$ $M^{*}(n,\epsilon,\sigma,\delta)\leq\sigma^{2}(n\delta\log e)^{1/2}$
$+$ $\sigma(2)^{1/2}(n\delta)^{1/4}(\log e)^{3/4}\cdot\Phi^{-1}(\epsilon)+O(\log n).$ (20)
The corresponding number of key bits should satisfy
$\log K(n,\epsilon,\sigma,\delta)=\begin{cases}\frac{1-\mu^{2}\sigma^{2}}{\mu}(\delta n\log e)^{1/2}+O(n^{1/4})\\ \,\ \ \ \ \ \ \ \ \ \ \ \,\text{if}\,\ \,\sigma^{2}<\frac{1}{\mu^{2}},\\ O(1)\,\ \ \ \ \ \,\text{otherwise}.\end{cases}$ (21)
In (1) and (21), $\mu\in(\frac{1}{2},1)$ is a hyper-parameter.
There are mainly four steps to arrive at achievability in Theorem 1.
  1. 1.
    Random coding scheme, decoding method, which are introduced in Section III-A. A code family consists of $MK$ codewords $\bm{x}_{sw},s=1,\cdots,K,w=1,\cdots,M$ are randomly generated according to truncated Gaussian distribution, and threshold decoding is adopted.
  2. 2.
    The reliability analysis (also in Section III-A) focuses on bounding two types of decoding error probability $\varepsilon_{sw}^{(1)}$ and $\varepsilon_{sw}^{(1)}$. The maximum of the first type is bounded in Lemma 4. The maximum of the second type can not be directly bounded. The expectation of it is bounded in Lemma 6, which leaves the bounding the maximum of the latter type for follow-up analysis.
  3. 3.
    The resolvability analysis in Section III-B focuses on bounding $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$. It also can not be directly bounded, and we provide an upper bound of its expectation in Theorem 5. Lemma 7, Lemma 8 and Lemma 9 are necessary preliminary results for Theorem 5.
  4. 4.
    The existence and characterization of the code family are provided in Section III-C. The expurgation operation is adopted to bound the maximum of the second type of error for each sub-code, meanwhile, the quantity $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ for the whole remaining codewords is shown to be low. The proof relies on the application of McDiarmid’s Inequality on codewords over AWGN channels.
There are two steps to arrive at converse in Theorem 1.
  1. 1.
    We show the converse constraint leads to an upper bound for the average power of the code family in Lemma 13.
  2. 2.
    By using the pigeon-hole principle, the converse bound for each sub-code under maximal power constraint is derived in Corollary 3, which further leads to characterization of the converse bound in Theorem 1.
For covert communication over AWGN channels, previous works [10] [11] have established the characterization of the first order asymptotics of the throughput as $O(\sqrt{n})$. This work focuses on refining these results by characterizing both the first and second order asymptotics as functions of the blocklength $n$, decoding error probability $\epsilon$, the background noise level parameter $\sigma$ and covert parameter $\delta$, which is not reported in [10] [11]. The analytic investigation stresses the influence of the ratio of the noise level (which is $\sigma^{2}/1$ in this work) on the throughput and the size of key bits. From Theorem 1, we can see that both of the first order asymptotics are $O(n^{\frac{1}{2}})$ and both of the second order asymptotic are $O(n^{\frac{1}{4}})$, and the only difference on the analytic coefficients is the hyper-parameter $\mu$ of truncation in the achievability. The truncation parameter $\mu$ lies in both the size of codebooks (1) and the size of the key (21), which is different from the analysis of the first order asymptotics of covert communication over AWGN channels [10] [11] [12]. The detailed explanation on our results will be provided in Section III and Section IV.
II-E Related Distributions
The following distributions will be important tool for us to understand how to construct random codes for covert communication over AWGN channels.
Definition 4.
  • Gaussian distributions array $P_{l}^{(m)}\sim\mathcal{N}(\bm{0},\mu\Psi(m)\bm{I}_{l})$ on $\mathbb{R}^{l}$ with $0<\mu<1$ where $\Psi(m)$ is the variance scaling function. The density function of $P_{l}^{(m)}$ is written as
    $g^{m}_{l}(\bm{x})=\frac{1}{(2\pi\mu\Psi(m))^{l/2}}e^{-\frac{\|\bm{x}\|^{2}}{2\mu\Psi(m)}}$ (22)
    or
    $g^{m}_{l}(x_{1},\cdots,x_{l})=\prod_{i=1}^{l}\frac{1}{\sqrt{2\pi\mu\Psi(m)}}e^{-\frac{x_{i}^{2}}{2\mu\Psi(m)}}.$ (23)
  • The truncated Gaussian distribution $\bar{P}_{l}^{(m)}$ with density function
    $\bar{g}^{m}_{l}(\bm{x})=\left\{\begin{split}&\frac{1}{\Delta}g^{m}_{l}(\bm{x}),\sqrt{\mu^{2}l\Psi(m)}\leq\hskip-2.8903pt\|\bm{x}\|\hskip-2.8903pt\leq\hskip-2.8903pt\sqrt{l\Psi(m)},\\ &0,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,otherwise\end{split}\right$. (24)
    and $\Delta_{l}^{m}$ is the normalized coefficient and equals to
    $\Delta_{l}^{m}=E_{P_{l}^{(m)}}[1_{\{\bm{x}\in\mathfrak{B}^{l}_{0}(\sqrt{l\Psi(m)})\backslash\mathfrak{B}^{l}_{0}(\sqrt{\mu^{2}l\Psi(m)})\}}],$ (25)
    where $\mathfrak{B}^{l}_{0}(r)$ is the $l$-dimensional sphere with center $\bm{0}$ and radius $r$.
  • Dirac measures. $1$-dimensional Dirac measure is defined as
    $\delta_{0}^{1}(x)=\left\{\begin{split}&+\infty,\ \ \ \ \ x=0\\ &0,\ \ \ \ \ \ \ \ \ \ x\neq 0\end{split}\right$. (26)
    and $l$-dimensional Dirac measure is defined as
    $\delta_{0}^{l}(\bm{x})=\prod_{i=1}^{l}\delta_{0}^{1}(x_{i}).$ (27)
In the definitions of $P_{l}^{(m)}$ and $\bar{P}_{l}^{(m)}$, $l$ denotes the dimension and $m$ denotes the variance scaling law. The key element in the design coding scheme for covert communications over AWGN channels lies in the how to choose the generating distribution of the random codes. In this work, we will show that Gaussian distributions with proper chosen parameters and their truncated versions lie in the core of both the achievability and converse parts. In the achievability, it depends on how to jointly design proper variance scaling function $\Psi$ under different blocklength $n$. The principle of the our design depends on the understanding of the distributions arrays involved with $\bar{P}_{l}^{(m)}$ and the Dirac measures $\delta_{0}^{l}(\bm{x})$.
For covert communication over AWGN channels, the following distributions are closely related. We introduce them according to where they lie at.
  • Distributions at the transmitter
    1. 1.
      Gaussian distribution $P_{n}^{(n)}$, which has probability density function $g^{n}_{n}(\bm{x})$ from (22) by letting $l=m=n$, we write as $g^{(n)}(\bm{x})$ for convenience.
    2. 2.
      The truncated Gaussian distribution $\bar{P}_{n}^{(n)}$ with density function
      $\bar{g}^{(n)}(\bm{x})=\left\{\begin{split}&\frac{1}{\Delta}g^{(n)}(\bm{x}),\sqrt{\mu^{2}n\Psi(n)}\leq\hskip-2.8903pt\|\bm{x}\|\hskip-2.8903pt\leq\hskip-2.8903pt\sqrt{n\Psi(n)},\\ &0,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,otherwise\end{split}\right$. (28)
      with $0<\mu<1$. The variable $\Delta$ is the normalized coefficient and equals to
      $\Delta=E_{P_{n}^{(n)}}[1_{\{\bm{x}\in\mathfrak{B}^{n}_{0}\big(\sqrt{n\Psi(n)}\big)\backslash\mathfrak{B}^{n}_{0}\big(\sqrt{\mu^{2}n\Psi(n)}\big)\}}],$ (29)
      where $\mathfrak{B}^{n}_{0}(r)$ is the $n$-dimensional sphere with center $\bm{0}$ and radius $r$.
  • Distributions at the legitimate receiver Bob.
    1. 1.
      The distribution of channel noise at the main channel with blocklength $n$: $Q^{(n)}_{\bm{Y},0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$, the probability density function is denoted as $f_{0}^{(n)}(\bm{y})$.
    2. 2.
      The output distribution $Q^{(n)}_{\bm{Y},1}\sim\mathcal{N}(\bm{0},(1+\mu\Psi(n))\bm{I}_{n})$ of the main channel corresponding to the input distribution $P_{n}^{(n)}$. The probability density function of $Q^{(n)}_{\bm{Y},1}$ is denoted as $f_{1}^{(n)}(\bm{y})$. From the definition of the channel, $f_{1}^{(n)}=g^{(n)}*f_{0}^{(n)}$.
    3. 3.
      The output distribution $\bar{Q}^{(n)}_{\bm{Y},1}$ of the main channel corresponding to the input distribution $\bar{P}_{n}^{(n)}$. The probability density function of $\bar{Q}^{(n)}_{\bm{Y},1}$ is denoted as $\bar{g}_{1}^{(n)}(\bm{y})$. From the definition of the channel, we have $\bar{f}^{(n)}_{1}=\bar{g}^{(n)}*f_{0}^{(n)}$.
  • Distributions at the adversary Willie.
    1. 1.
      The distribution of channel noise at the adversary with blocklength $n$: $Q^{(n)}_{\bm{Z},0}(\bm{z})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}_{n})$, the probability density function of which is denoted as $h_{0}^{(n)}(\bm{z})$.
    2. 2.
      The mixture Gaussian distribution induced by the codebook at the adversary $Q^{(n)}_{\bm{Z},\bm{C}}$, which is defined in (13).
    3. 3.
      The output distribution $Q^{(n)}_{\bm{Z},1}$ at the adversary corresponding to the input distribution $P_{n}^{(n)}$. The probability density function of $Q^{(n)}_{\bm{Z},1}$ is denoted as $h_{1}^{(n)}(\bm{z})$. From the definition of the channel, we have $h^{(n)}_{1}=g^{(n)}*h_{0}^{(n)}$.
    4. 4.
      The output distribution $\bar{Q}^{(n)}_{\bm{Z},1}$ of the channel corresponding to the input distribution $\bar{P}_{n}^{(n)}$. The probability density function of $\bar{Q}^{(n)}_{\bm{Z},1}$ is denoted as $\bar{h}_{1}^{(n)}(\bm{z})$. From the definition of the channel, we have $\bar{h}_{1}^{(n)}=\bar{g}^{(n)}*h_{0}^{(n)}$.
Note that the distributions $P_{n}^{(n)}$ and $\bar{P}_{n}^{(n)}$ are the special cases of Gaussian distributions array and their truncated version in previous definition. The next lemma and two corollaries characterize the closeness of the distributions $P_{n}^{(n)}$ and $\bar{P}_{n}^{(n)}$.
Lemma 2.
The KL divergence and TVD between $\bar{P}_{n}^{(n)}$ and $P_{n}^{(n)}$ are
$\mathbb{D}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=-\log\Delta,$ (30)
$\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=1-\Delta.$ (31)
with
$\Delta=\frac{\gamma(n/2,n/2\mu)-\gamma(n/2,n\mu/2)}{\Gamma(n/2)},$ (32)
where $\gamma(a,z)$ is incomplete gamma function defined as follows
$\gamma(a,z)=\int_{0}^{z}e^{-t}t^{a-1}dt.$ (33)
The proof can be found in Appendix B. With $0<\mu<1$ being a constant, $\Delta\rightarrow 1$ as $n\rightarrow\infty$ exponentially and the rapidity depends only on $\mu$. Please see Fig.2 for reference. Intuitively, the truncation of the tail of the Gaussian family depends on the hyper-parameter $\mu$. When $\mu$ is close to $1$, the tail probability is exponentially decreasing. Thus, we have the following corollary for $\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})$, the details of the analysis has been presented in Section IV-A of [23].
Fig. 2: $\Delta$ tends to 11 with proper chosen $\mu$, which is the numerical results from (32).
Corollary 1.
For proper fixed $\mu$, we have
$\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})\rightarrow 0\,\ \ \,\text{as}\,\ \ \,n\rightarrow\infty.$ (34)
Let $C_{0}(-\infty,\infty)$ be the class of bounded continuous functions vanishing at $\infty$ and $C[-\infty,\infty]$ be the class of continuous functions whose limits at -$\infty$ and $\infty$ are finite, which are both the subclasses of bounded continuous functions and have relationship as $C_{0}(-\infty,\infty)\subset C[-\infty,\infty]$. We first recall the following theorem (Theorem 1 at page 249 of [36]).
Theorem 2.
In order that a probability distribution $F_{n}$ converges properly [^1] Proper convergence is equivalent to weak convergence with respect to $C_{0}(-\infty,\infty)$. to $F$, it is necessary and sufficient that
$\mathbb{E}_{n}(u)\rightarrow\mathbb{E}(u)\,\ \ \,\text{for all}\,\ \ u\in C_{0}(-\infty,\infty)$ (35)
where $\mathbb{E}_{n}$ and $\mathbb{E}$ are expectation with respect to $F_{n}$ and $F$.
Let $\|\cdot\|_{m}^{*}$ be a metric of proper convergence of probability measures, we have the following corollary.
Corollary 2.
For proper fixed $\mu$, the distribution pair $\bar{P}_{n}^{(n)}$ and $P_{n}^{(n)}$ satisfies
$\|\bar{P}_{n}^{(n)}-P_{n}^{(n)}\|_{m}^{*}\rightarrow 0.$ (36)
Proof.
For any continuous function $u(\bm{x})$ on $\mathbb{R}^{n}$ which vanishes at infinity, we have
$\begin{split}&|\int u(\bm{x})P_{n}^{n}(d\bm{x})-\int u(\bm{x})\bar{P}_{n}^{n}(d\bm{x})|\\ \leq&\int u(\bm{x})|g^{n}(\bm{x})-\bar{g}^{n}(\bm{x})|d\bm{x}\\ \leq&\,\,2\,\,\underset{\bm{x}}{\sup}|u(\bm{x})|\cdot\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)}).\end{split}$
From the assumption of $u$ and Corollary 1, we have
$\mathbb{E}_{\bar{P}_{n}^{(n)}}(u)\rightarrow\mathbb{E}_{P_{n}^{(n)}}(u).$ (37)
By using a multivariate version of Theorem 2, we have the conclusion. ∎
Different from discrete memoryless channels, the input distribution and the output distributions of AWGN channels are related by the convolution operation with the distribution of channel noise. In the following analysis, we treat the input distributions as convolution operators, and leverage the metric of proper convergence in probability theory to characterize the closeness of the output distributions at both the main channel $W_{1}$ and the channel at the adversary $W_{2}$. The following definition is necessary, which can be found in [36].
Definition 5.
Let $F$ be a probability distribution and $\phi$ a bounded function, the convolution of $\phi$ with $F$ is the resulting function $u$ as
$u(\bm{y})=\int_{x_{n}}\cdots\int_{x_{1}}\phi(\bm{y}-\bm{x})F(d\bm{x})$ (38)
Following the language of operators, it will be denoted as $u=F\phi$. When $F$ has a density function $f$, it is written alternatively $u=f*\phi$.
The following theorem from [36] (page 255) is necessary for our discussion.
Theorem 3.
A sequence of probability distributions $P_{n}$ converges properly to a probability distribution $P_{0}$ iff for each $u\in C[-\infty,\infty]$, the convolutions $U_{n}=\mathfrak{F}_{n}u$ converge uniformly to a limit $U=\mathfrak{F}u$.
For the terms $\frac{\bar{Q}^{(n)}_{\bm{Y},1}(\bm{y})}{Q^{(n)}_{\bm{Y},1}(\bm{y})}$ and $\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}$, we have the following bounds when $n$ is sufficiently large.
Lemma 3.
For sufficiently large $n$, for any proper chosen $\mu>0.5$ and any $\bm{z}$, we have
$1-\iota\leq\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\leq 1+\iota,$ (39)
$1-\iota^{\prime}\leq\frac{\bar{h}_{1}^{(n)}(\bm{z})}{h^{(n)}_{1}(\bm{z})}\leq 1+\iota^{\prime}.$ (40)
where $\iota,\iota^{\prime}\rightarrow 0$ as $n\rightarrow\infty$.
The proof can be found in Appendix C and is an application of Theorem 3 with the specific distributions $P_{n}^{(n)}$ and $\bar{P}_{n}^{(n)}$.
Remark 4.
In the follow-up analysis, we don’t distinguish $\bar{Q}^{(n)}_{\bm{Y},1}(\bm{y})$, $Q^{(n)}_{\bm{Y},1}(\bm{y})$, $\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})$ and $Q^{(n)}_{\bm{Z},1}(\bm{z})$ from their probability density functions. These expressions make our results consistent with that of DMCs and avoid utilization of too many symbols. In the following sections, we will continue to adopt these expressions for convenience.
III Throughput Analysis of Covert Communications Over AWGN channels
This section provides main results - the first and second order asymptotics of the maximal throughput of covert communication over AWGN channels. The results are organized as follows. In Section III-A, we present the random coding scheme, the decoding method and the analysis of decoding error probability. In Section III-B, channel resolvability results are developed.
III-A Random Coding, Decoding and Reliability Analysis
This section present one-shot achievability results. This part differs from previous works over DMCs on the choices of the generating distribution and the reference distribution for decoding. The proofs follow the same line as [22] by considering the two types of error $\varepsilon_{sw}^{(1)}$ and $\varepsilon_{sw}^{(2)}$ separately. Nevertheless, as the generating distribution are truncated Gaussian, whose support set is a subset of $\mathbb{R}^{n}$, the error probability $\varepsilon_{sw}^{(1)}$ can be considered separately for codewords with different radius in $\mathbb{R}^{n}$. By properly choosing the decoding threshold, we ensure that $\varepsilon_{sw}^{(1)}$ for each $(s,w)$ pair is small. For error in the second type $\varepsilon_{sw}^{(2)}$, we bound the expectation of it for the generation distribution, and we will use expurgation operation later to pick part of codewords such that this type of errors of them are also small.
The transmitter Alice generates a family of $K$ codebooks $\{\bm{x}_{sw},s\in\langle K\rangle,w\in\langle M\rangle\}$ from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$ by independent and identically distributed sampling. Then, she choose an index $s$ from $\langle K\rangle$. The unique key $\mathsf{S}(s)$ is sent to the legitimate receiver Bob and is unknown to the adversary. With the key $\mathsf{S}(s)$ , Alice choose the $s$-th codebook to communicate with Bob. Meanwhile, Bob utilizes the $s$-th codebook in the threshold decoding upon observing a channel output vector $\bm{y}$. With a parameter $\gamma$ and a fixed distribution $Q^{(n)}_{\bm{Y},1}$ on $\mathcal{R}^{n}$, the decoder decides $\widehat{W}=w$ as the transmitted message if $w$ is the unique index in $\langle M\rangle$ such that
$\log\frac{W_{1}^{n}(\bm{y}|{\bm{x}_{sw}})}{Q^{(n)}_{\bm{Y},1}(\bm{y})}>\gamma.$ (41)
If there is no $w$ satisfying (41), the decoder decares an error. This decoder has been adopted in the one shot achievability analysis of [22] and is based on the threshold decoder [20].
The conditional probability of error when $W=w$ is upper bounded by the sum of the following two types.
[…]
$\varepsilon_{sw}^{(2)}=\mathbb{P}\left(\exists{w}^{\prime}\neq w:\log\frac{W_{1}^{n}(\bm{Y}|{\bm{x}_{sw^{\prime}}})}{Q^{(n)}_{\bm{Y},1}(\bm{Y})}>\gamma\right).$ (43)
Note that $Q^{(n)}_{\bm{Y},1}$ is the induced distribution of the channel corresponding to $P_{n}^{(n)}$, which is different from the induced distribution $\bar{Q}^{(n)}_{\bm{Y},1}$ of the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$.
As maximal error probability is the reliability metric, we consider the two types of error $\varepsilon_{sw}^{(1)}$ and $\varepsilon_{sw}^{(2)}$ separately. The bounding of the first type of error will make use of the Berry-Esseen Theorem, which we recall here for convenience.
Theorem 4.
(Berry-Esseen Theorem): Let $X_{1},\cdots,X_{n}$ be independent random variables such that for $i\in\langle n\rangle$, $\mathbb{E}(X_{i})=u_{i}$, $Var(X_{i})=v_{i}^{2}$ and $t_{i}=\mathbb{E}\big(|X_{i}-\mu_{i}|^{3}\big)$. If we define $v^{2}=\sum_{i=1}^{n}v_{i}^{2}$ and $t=\sum_{i=1}^{n}t_{i}$, then we have
$|\mathbb{P}\bigg(\sum_{i=1}^{n}(X_{i}-u_{i})\leq\lambda v\bigg)-\Phi(\lambda)|\leq\frac{6t}{v^{3}}.$ (44)
The following lemma bounds the maximum of $\varepsilon_{sw}^{(1)}$ by choosing proper decoding threshold $\gamma$. As the power of the codewords ranges from $\mu^{2}n\Psi(n)$ to $n\Psi(n)$, the bounding of the error is slightly different from previous works [20][24], where sphere symmetry can be leveraged for codewords with equal power. Nevertheless, by separately considering the codewords with equal power and choosing an universal decoding threshold, the analysis can also be simplified by leveraging sphere symmetry.
Lemma 4.
For any given $\epsilon_{1}>0$, if we choose
$\gamma=$ $nC(\mu,\Psi(n))+n\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))$
$+$ $\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}$
$\times$ $\Phi^{-1}\bigg(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\bigg)$ (45)
where
$\mathsf{C}(\mu,\Psi(n))=\frac{1}{2}\log(1+\mu\Psi(n)),$ (46)
$\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))=\frac{\mu^{2}\Psi(n)-\mu\Psi(n)}{2(1+\mu\Psi(n))}\log e,$ (47)
$\mathsf{V}(\Psi(n),\mu\Psi(n))=\frac{2\Psi(n)+\mu^{2}\Psi^{2}(n)}{2(1+\mu\Psi^{2}(n))}\log^{2}e,$ (48)
and $\Phi$ is the cumulative distribution function of standard normal random variable, then we have
$\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(1)}\leq\epsilon_{1}.$ (49)
Proof.
For a codeword $\bm{x}_{sw}$ being sampled from truncation Gaussian distribution $\bar{P}_{n}^{(n)}$, let $\|\bm{x}_{sw}\|^{2}=n\Upsilon$, where $\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$. We consider the probability of error in (42) for $\bm{x}_{sw}$ on the n1n-1 dimensional sphere with radius $\sqrt{n\Upsilon}$. Due to the sphere symmetry of the reference distribution $Q^{(n)}_{\bm{Y},1}$, we assume $\bm{x}_{0}=[\sqrt{\Upsilon},\cdots,\sqrt{\Upsilon}]$. The sphere symmetry has been employed in previous works [24] and [20] to simplify the analysis of decoding error. The information density $\log\frac{W_{1}^{n}(\bm{Y}|{\bm{x}_{0}})}{Q^{(n)}_{\bm{Y},1}(\bm{Y})}=\log\frac{W_{1}^{n}(\bm{Y}|{\bm{x}_{0}})}{f_{1}^{(n)}(\bm{y})}$ is expressed as
$\frac{n}{2}\log\big(1+\mu\Psi(n)\big)+\frac{\log e}{2}\sum_{i=1}^{n}\left[\frac{y_{i}^{2}}{1+\mu\Psi(n)}-(y_{i}-\sqrt{\Upsilon})^{2}\right].$ (50)
Under $W_{1}^{n}(\bm{Y}|{\bm{x}_{0}})$, it has the same distribution as
$\begin{split}H^{\Upsilon}_{n}=&\frac{n}{2}\log\big(1+\mu\Psi(n)\big)\\ +&\frac{\log e}{2(1+\mu\Psi(n))}\sum_{i=1}^{n}\left[\Upsilon-\mu\Psi(n)\xi_{i}^{2}+2\sqrt{\Upsilon}\xi_{i}\right]\end{split}$ (51)
where we have used $\xi_{i}\sim\mathcal{N}(0,1)$ for $i=1,\cdots,n$. We rewrite (51) as $n\mathsf{C}(\mu,\Psi(n))+\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n))$, where $\mathsf{C}(\mu,\Psi(n))$ is defined in (46) and
$S_{i}(\Upsilon,\mu\Psi(n))=\frac{\log e}{2(1+\mu\Psi(n))}\left(\Upsilon-\mu\Psi(n)\xi_{i}^{2}+2\sqrt{\Upsilon}\xi_{i}\right).$ (52)
As $S_{i}(\Upsilon,\mu\Psi(n)),i=1,\cdots,n$ are i.i.d random variables, we have
$\begin{split}\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n))\right]=&\frac{\Upsilon-\mu\Psi(n)}{2(1+\mu\Psi(n))}\log e\\ =&\bar{\mathsf{C}}\big(\Upsilon,\mu\Psi(n)\big),\end{split}$ (53)
$\begin{split}\mathsf{Var}\left[\frac{1}{n}\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n))\right]=&\frac{1}{n}\frac{2\Upsilon+\mu^{2}\Psi^{2}}{2(1+\mu\Psi)^{2}}\log^{2}e\\ =&\frac{\mathsf{V}\big(\Upsilon,\mu\Psi(n)\big)}{n},\end{split}$ (54)
$\mathsf{T}(\Upsilon,\Psi(n)):=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[|S_{i}(\Upsilon,\mu\Psi(n))-\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))|^{3}\right].$ (55)
From the above (53) and (54), we further have $\forall\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$, the following inequalities hold,
$\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))\leq\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))\leq\bar{\mathsf{C}}(\Psi(n),\mu\Psi(n)),$ (56)
$\frac{\mathsf{V}(\Psi(n),\mu\Psi(n))}{\mathsf{V}(\Upsilon,\mu\Psi(n))}=\frac{2\Psi(n)+\mu^{2}\Psi^{2}}{2\Upsilon+\mu^{2}\Psi^{2}}\geq 1,$ (57)
In addition, we have
$\bar{\mathsf{C}}(\mu\Psi(n),\mu\Psi(n))=0,$ (58)
and
$\mathsf{V}(\mu\Psi(n)):=$ $\mathsf{V}(\mu\Psi(n),\mu\Psi(n))$
$=$ $\frac{2\mu\Psi(n)+\mu^{2}\Psi^{2}(n)}{2(1+\mu\Psi(n))^{2}}.$ (59)
Let $B(\Upsilon,\mu\Psi(n))=\frac{6\mathsf{T}\left[S_{i}(\Upsilon,\mu\Psi(n))\right]}{V[S_{i}(\Upsilon,\mu\Psi(n))]^{\frac{3}{2}}}$ and it is easy to obtain that $B(\Upsilon,\mu\Psi(n))$ is bounded for any $\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$. Hence, for any given $0<\epsilon_{1}<1$, if $n$ is sufficiently large, we have
$\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}>0.$ (60)
Now, let $\gamma$ be as in (4). Obviously, $\gamma$ does not depend on $\Upsilon$. We rewrite the term $\mathbb{P}_{W_{1}^{n}(\bm{Y}|\bm{X}=\bm{x}_{0})}\left(\log\frac{W_{1}^{n}(\bm{Y}|{\bm{X}})}{{g}_{1}^{(n)}(\bm{Y})}<\gamma\right)$ in (III-A).
$\mathbb{P}_{W_{1}^{n}(\bm{Y}|\bm{X}=\bm{x}_{0})}\left(\log\frac{W_{1}^{n}(\bm{Y}|{\bm{X}})}{{g}_{1}^{(n)}(\bm{Y})}<\gamma\right)$
$=$ $\mathbb{P}\left(\sum_{i=1}^{n}S_{i}(\Upsilon,\Psi(n))<\gamma-n\mathsf{C}(\mu,\Psi(n))\right)$
$\overset{(a)}{\leq}$ $\Phi\left(\frac{\gamma-n\mathsf{C}(\mu,\Psi(n))-n\cdot\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))}{\sqrt{n\mathsf{V}(\Upsilon,\Psi(n))}}\right)+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$=$ $\Phi\left(\frac{n\cdot\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))-n\cdot\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))+\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}\Phi^{-1}\big(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\big)}{\sqrt{n\mathsf{V}(\Upsilon,\Psi(n))}}\right)$
$+$ $\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$\overset{(b)}{\leq}$ $\Phi\left(\sqrt{\frac{\mathsf{V}(\Psi(n),\mu\Psi(n))}{\mathsf{V}(\Upsilon,\Psi(n))}}\Phi^{-1}\bigg(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\bigg)\right)+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$\overset{(c)}{\leq}$ $\Phi\left(\Phi^{-1}\bigg(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\bigg)\right)+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$=$ $\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}+\frac{B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}$
$\leq$ $\epsilon_{1}$ (61)
In the derivation of (III-A), the step (a) follows from applying Berry-Esseen Theorem to i.i.d random variables $S_{i}(\Upsilon,\mu\Psi(n)),i=1,\cdots,n$, the step (b) follows from (56): $\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))-\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))\leq 0$, and step (c) follows from that $\Phi^{-1}\big(\epsilon_{1}-\frac{\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}}\big)<0$ and (57): $\frac{\mathsf{V}(\Psi(n),\mu\Psi(n))}{\mathsf{V}(\Upsilon,\Psi(n))}\geq 1$. From the arguments above, the bound is applicable with any $\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$. Due to the sphere symmetry, we have for any randomly generated codeword $\bm{x}_{s,w}$ from $\bar{P}_{n}^{(n)}$, $\varepsilon_{s,w}^{(1)}<\epsilon_{1}$, i.e.,
$\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(1)}\leq\epsilon_{1}.$ (62)
Remark 5.
In the $\kappa\beta$ bound (Theorem 25) of [20], the codewords lie on the $n$-dimensional sphere $\mathcal{S}^{n-1}$ with radius $\sqrt{nP}$ and are not randomly generated by any distribution. Nevertheless, the auxiliary distribution $\mathcal{N}(0,(1+P)\bm{I}_{n})$ is chosen so that normal approximation can be performed. In this work, the codewords lie on the gray region as in Fig. 3.
Fig. 3: The codewords lie in the gray region.
The generating distribution $\bar{P}_{n}^{(n)}$ is chosen such that its induced output distribution $\bar{Q}^{(n)}_{\bm{Y},1}$ is close to the auxiliary distribution $Q^{(n)}_{\bm{Y},1}$, which makes the normal approximation accurate. The strategy of bounding the error probability of the first type here is separately considering $n$-dimensional sphere $\mathcal{S}^{n-1}$ with radius $\sqrt{n\Upsilon}$, which ranges from $\sqrt{\mu^{2}n\Psi(n)}$ to $\sqrt{n\Psi(n)}$. For each sphere, we can utilize the sphere symmetry to simplify the analysis. The parameter $\mu$ makes the interval under control so that an universal decoding threshold $\gamma$ can be obtained. As a consequence, there is an additional term $\bar{\mathsf{C}}(\Upsilon,\mu\Psi(n))$ in the normal approximation, which is different from the previous works [24][20] where all the codewords have the same power level. The bounding of the probability of error $\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(1)}$ relies on the following facts. As $\mu$ is close to $1$, the $l^{2}$ norms of the codewords lie in a small interval $\mu^{2}n\Psi(n),n\Psi(n)]$. If we choose a small $\mu$ (close to 0), the size of the codebook should also be small so as to bound the error probability of the first type. This will not emerge when we only consider the first order asymptotics.
Remark 6.
In the evaluation of RCU bound in [24], the codewords are uniformly distributed on $\mathcal{S}^{n-1}$ with radius $\sqrt{nP}$, and maximal likelihood decoding is adopted. The derivation relies on the high-probability set $\mathcal{F}$ (Please refer to (31) in Section IV-D of [24]) and the fact that the induced output distribution and the auxiliary distribution can be bounded by a finite constant:
$\underset{\bm{y}\in\mathcal{F}}{\sup}\frac{f_{\bm{X}W^{n}}(\bm{y})}{f^{*}_{\bm{Y}}(\bm{y})}\leq J.$ (63)
In our setting, the above techniques can not be applied. A similar high-probability set here is expressed as
$\mathcal{F}=$ $\left\{\bm{y}\in\mathbb{R}^{n}:\frac{1}{n}\times\|\bm{y}\|^{2}\in\right$.
$\left.\big[1+\mu^{2}\Psi(n)-\tau_{n},1+\Psi(n)+\tau_{n}\big]\right\}.$ (64)
In this case, the inequality like (63) no longer holds when $f^{*}_{\bm{Y}}$ is the pdf of the particular auxiliary distribution $Q^{(n)}_{\bm{Y},1}$.
The last lemma has provided a bound on error probability of the first type for all candidate codewords as: $\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(1)}\leq\epsilon_{1}$ for given $\epsilon_{1}$. While, for the error probability of the second type $\varepsilon_{sw}^{(2)}$, we can not bound it for each codeword. Nevertheless, our strategy is divided into two steps. First, we use an existing results to bound the average of it in Lemma 6. Then, we leverage the expurgation later to discard part of codewords with large error probability of the second type. The following lemma is almost a restatement of part of Lemma 1 in [22]. The differences are: (1) The discrete memoryless channel is substituted by the AWGN channel. (2) The sum is substituted by integration. (3) We use $\log$ with base $2$ instead of base $e$.
Lemma 5.
For the AWGN channel (5), let $P_{X^{n}}$, $P_{Y^{n}}$ and $Q_{Y^{n}}$ be distributions on $\mathbb{R}^{n}$, where $P_{Y^{n}}$ is the induced distribution of $P_{X^{n}}$. If we choose $\bm{x}_{1},\cdots,\bm{x}_{M}$ independently according to $P_{X^{n}}$, then by the decoding rule (41), for all $\gamma$ and $w\in\langle M\rangle$
$\mathbb{E}\left(\varepsilon_{sw}^{(2)}\right)=M\cdot 2^{-\gamma}\mathbb{E}_{P_{Y^{n}}}\bigg(\frac{P_{Y^{n}}}{Q_{Y^{n}}}\bigg).$ (65)
Applying Lemma 5 for (43) in our setting, we have
$\begin{split}\mathbb{E}\left(\epsilon_{sw}^{(2)}\right)=&M_{n}\cdot 2^{-\gamma}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{Q}^{(n)}_{\bm{Y},1}}{Q^{(n)}_{\bm{Y},1}}\bigg)\\ =&M_{n}\cdot 2^{-\gamma}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\end{split}$ (66)
The above upper bound on the average of the error probability of the second type can be further elaborated if we choose a proper parameter $M$.
Lemma 6.
If we choose
$\begin{split}M=M_{n}=e^{\gamma-c_{0}\log n},\end{split}$ (67)
for any positive constant $c_{0}$, then
[…]
and
$\log M_{n}=$ $nC(\mu,\Psi(n))+n\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))$
$+$ $\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}$
$\times$ $\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})$
- $c_{0}\log n.$ (69)
The proof is obvious from (66) and (4). Note that here we just provide an upper bound of the average error probability of the second type $\varepsilon_{sw}^{(2)}$. To get a reliable code, we will adopt expurgation later in Theorem 7 to construct a reliable code such that the maximum of the sum $\varepsilon_{sw}^{(1)}+\varepsilon_{sw}^{(2)}$ for all $(s,w)$ pairs is small.
Remark 7.
From (68), the average of the decoding error probability of the second type depends on two elements. The first element is $\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$. The integrand is the ratio of the probability density functions of the induced output distribution of the generating distribution and the reference distribution, where the latter is the induced distribution of Gaussian distribution without truncation. The second element is the size of the codebook, which has been almost determined in Lemma 4. Hence, it is necessary to control the effect of truncation to bound the ratio $\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}$, which depends on the choice of the parameter $\mu$ and has been uniformly bounded for sufficiently large $n$ in Lemma 3.
III-B Resolvability Analysis
This section presents one-shot resolvability analysis of covert communication over AWGN channels, i.e., the divergence between the induced distribution of the codebook and the distribution of noise: $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$. Different from the binary input discrete channels in [22], where the resolvability analysis resorts to modified triangle inequality of divergences as quasi-metrics, the quasi-metric property of KL divergence no longer holds over the AWGN channel because here the involved distributions lie on $\mathbb{R}^{n}$.
Consider the codewords $\bm{X}_{sw},s\in\langle K\rangle,w\in\langle M\rangle$ as random variables. The KL divergence $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ is expressed as
$\begin{split}&\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{z},0})\\ =&\int_{z}\sum_{s=1}^{K}\sum_{w=1}^{M}\frac{1}{MK}W^{n}_{2}(\bm{z}|\bm{X}_{sw})\\ \times&\log\frac{\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{X}_{sw})}{MKQ^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z},\end{split}$ (70)
which is difficult to analyze directly. We rewrite it as
$\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$
$=\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\left[\log\frac{Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})}{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}+\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}\right$.
$+\left.\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}\right]d\bm{z}$
$=$ $\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})}{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}$
$+$ $\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\cdot\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}$
$+$ $\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$ (71)
Note that the three terms in (III-B) are all random variables as they are functions of $\bm{X}_{sw},s=1,\cdots,K,w=1,\cdots,M$, which are randomly generated from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$. Our strategy of bounding $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ is as follows. We directly investigate the expectation of each term in (III-B). By showing that the expectation of each term in (III-B) is small, it is reasonable that the event that the realization of $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ induced by a set of legitimate codewords is small has positive probability. From the linearity of the expectation, we have
$\begin{split}\mathbb{E}&\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\right]\\ =&\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|\bar{Q}^{(n)}_{\bm{Z},1})\right]+\mathbb{E}\left[\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\cdot\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}\right]\\ +&\mathbb{E}\left[\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}\right].\end{split}$ (72)
The follow-up analysis will focus on upper bounding the three terms in (72), among which a large body of analysis will be taken on the first term and an upper bound of it is provided in Lemma 9. Here we outline the idea of the proof. First, we apply the following inequality (the inequality (10) in [37]), which has also been used in Lemma 3 in [22] for resolvability analysis for DMCs.
$\begin{split}&\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|\bar{Q}^{(n)}_{\bm{Z},1})\right]\\ \leq&\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{X})\bar{P}_{n}^{(n)}(\bm{X})}\left(\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\right).\end{split}$ (73)
Then, to upper bound the term (73), we divide the product space $\bm{X}\times\bm{Z}$ into two parts. For the first part, we show that the inner term $\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)$ for each $(\bm{x},\bm{z})$ pair is under control. For the second part, we prove that the probability of it is exponentially decreasing. In fact, a precise upper bounding of the first part depends on the choice of the sizes of codebook and the key, which will be illustrated later in Theorem 7.
Before establishing Lemma 9, we first define $\mathcal{E}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}\leq\gamma_{1}\}$. Note that in the definition of $\mathcal{E}$, we use ${Q}^{(n)}_{\bm{Z},1}$ but not $\bar{Q}^{(n)}_{\bm{Z},1}$. Then, some technical preparations are provided in Lemma 7 and Lemma 8. Lemma 7 is on the ratio $\frac{W_{2}^{n}(\bm{z}|\bm{x})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}$, the proof of which can be found in Appendix D.
Lemma 7.
$\frac{W_{2}^{n}(\bm{z}|\bm{x})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}\leq\bigg(\frac{\sigma^{2}+\mu\Psi(n)}{\sigma^{2}}\bigg)^{\frac{n}{2}}e^{\frac{n}{2\mu}}$ (74)
For the event $\mathcal{E}^{c}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}>\gamma_{1}\}$, the following Lemma 8 provides an upper bound on its probability with proper chosen $\gamma_{1}$. Similar with the proof of Lemma 4, our strategy here is separately considering the codewords with equal power and choosing an universal threshold $\gamma_{1}$. As a consequence, the sphere symmetry can be leveraged to show that the probability of the event $\{\bm{z}:\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}>\gamma_{1}\}$ for each $\bm{x}$ has large-deviation bounds.
Lemma 8.
If we choose
$\gamma_{1}=n\mathsf{C}(\mu,\Psi,\sigma^{2})-n(1+\rho)\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})$ (75)
with any $\rho>0$, where
$\mathsf{C}(\mu,\Psi,\sigma^{2})=\frac{1}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})$ (76)
and
$\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})=\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}(\mu\Psi(n)-\Psi(n)),$ (77)
then
$\mathbb{P}(\mathcal{E}^{c})\leq ae^{-bn}$ (78)
with some $a,b>0$ as functions of $\rho$.
Proof.
The term $\mathbb{P}(\mathcal{E}^{c})$ can be rewritten as follows
$\begin{split}&\mathbb{P}(\mathcal{E}^{c})=\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})W^{n}_{2}(\bm{Z}|\bm{X})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|\bm{X})}{{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}>\gamma_{1}\bigg)\\ =&\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})}\left[\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)\right].\end{split}$ (79)
Note that in (79), the distribution ${Q}^{(n)}_{\bm{Z},1}$ is the induced distribution of the channel $W_{2}^{n}$ by $P_{n}^{(n)}$ but not $\bar{P}_{n}^{(n)}$, which is an isotropic Gaussian distribution. The inner term in (79)
$\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)$ (80)
is considered for each $\bm{x}_{sw}$ in the support set of $\bar{P}_{n}^{(n)}$. Similarly as the proof of Lemma 4, we first consider the probability of error in (80) for $\bm{x}_{sw}$ on the $n$ dimensional sphere with radius $\sqrt{n\Upsilon}$. The sphere symmetry allows us to assume $\bm{x}_{0}=[\sqrt{\Upsilon},\cdots,\sqrt{\Upsilon}]$. we assume a codeword $\bm{x}_{0}=[\sqrt{\Upsilon},\cdots,\sqrt{\Upsilon}]$, where $\Upsilon\in[\mu^{2}\Psi,\Psi]$. The information density $\log\frac{W_{2}^{n}(\bm{z}|{\bm{x}_{0}})}{h_{1}^{(n)}(\bm{z})}$ is expressed as
$\begin{split}&\frac{n}{2}\log\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)\\ +&\frac{\log e}{2}\sum_{i=1}^{n}\left[\frac{z_{i}^{2}}{\sigma^{2}+\mu\Psi(n)}-\frac{(z_{i}-\sqrt{\Upsilon})^{2}}{\sigma^{2}}\right]\\ =&\frac{n}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})-\sum_{i=1}^{n}\frac{\log e}{2\sigma^{2}(\sigma^{2}+\mu\Psi(n))}\\ \times&\left[\mu\Psi(n)z_{i}^{2}-2(\sigma^{2}+\mu\Psi(n))\sqrt{\Upsilon}z_{i}+(\sigma^{2}+\mu\Psi(n))\Upsilon\right]\end{split}$ (81)
Under $W_{2}^{n}(\bm{Z}|{\bm{x}_{0}})$, we have $z_{i}=\sqrt{\Upsilon}+\sigma\xi_{i}$, then (81) has the same distribution as
$\bar{H}^{\Upsilon}_{n}=$ $\frac{n}{2}\log\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)$
$+$ $\frac{\log e}{2\sigma^{2}(\sigma^{2}+\mu\Psi(n))}\sum_{i=1}^{n}\bigg(\sigma^{2}\Upsilon-\mu\Psi(n)\sigma^{2}\xi_{i}^{2}$
$+$ $2\sigma^{3}\sqrt{\Upsilon}\xi_{i}\bigg)$ (82)
where we have used $\xi_{i}\sim\mathcal{N}(0,1)$ for $i=1,\cdots,n$. We rewrite (III-B) as $n\ \mathsf{C}(\mu,\Psi(n),\sigma^{2})-\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})$, where $\mathsf{C}(\mu,\Psi(n),\sigma^{2})=\frac{1}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})$ and
$\begin{split}&S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})\\ =&\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}\left(\mu\Psi(n)\xi_{i}^{2}-\Upsilon-2\sigma\sqrt{\Upsilon}\xi_{i}\right)\\ =&\frac{\mu\Psi(n)\log e}{2(\sigma^{2}+\mu\Psi(n))}\left[(\xi_{i}-\frac{\sigma\sqrt{\Upsilon}}{\mu\Psi(n)})^{2}-\frac{(\sigma^{2}+\mu\Psi(n))\Upsilon}{\mu^{2}\Psi^{2}(n)}\right]\\ =&\frac{\mu\Psi(n)\log e}{2(\sigma^{2}+\mu\Psi(n))}\theta_{i}^{2}-\frac{\Upsilon\log e}{2\mu\Psi(n)},\end{split}$ (83)
with $\theta_{i}=\xi_{i}-\frac{\sigma\sqrt{\Upsilon}}{\mu\Psi(n)}$. It is obvious that $S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2}),i=1,\cdots,n$ are i.i.d random variables. For each $i$ and any $\Upsilon\in[\mu^{2}\Psi,\Psi]$
$\begin{split}\mathbb{E}\left[S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})\right]=&\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}(\mu\Psi(n)-\Upsilon)\\ =&\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2}).\end{split}$ (84)
Furthermore, we have $\forall\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$,
$\begin{split}\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\leq&\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})\\ \leq&\hat{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n),\sigma^{2}),\end{split}$ (85)
and
$\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})=\frac{\log e}{2(\sigma^{2}+\mu\Psi(n))}(\mu-1)\Psi(n)<0.$ (86)
Now, let $\gamma_{1}$ be as in (75) with $\rho>0$, which does not depend on $\Upsilon$. The inequality (80) can be rewritten as (III-B), where the step (a) follows from (83), and step (b) follows from the tail probability bound of chi square distribution in (Theorem 4 in [32]), (85) and (86). The step (b) can also be derived by applying Chernoff bounding technique for sub-exponential random variables [33].
$\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x}_{0})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}_{0}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)$
$=$ $\mathbb{P}\left[n\ \mathsf{C}(\mu,\Psi(n),\sigma^{2})-\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})>n\mathsf{C}(\mu,\Psi(n),\sigma^{2})-n(1+\rho)\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\right]$
$=$ $\mathbb{P}\left[\sum_{i=1}^{n}S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})<n(1+\rho)\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\right]$
$=$ $\mathbb{P}\left[\sum_{i=1}^{n}\left(S_{i}(\Upsilon,\mu\Psi(n),\sigma^{2})-\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})\right)<n\left(\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})-\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})\right)+n\rho\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\right]$
$\overset{(a)}{=}$ $\mathbb{P}\left[\frac{\mu\Psi(n)\log e}{2(\sigma^{2}+\mu\Psi(n))}\sum_{i=1}^{n}\left(\theta_{i}^{2}-E(\theta_{i}^{2})\right)<n\left(\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})-\hat{\mathsf{C}}(\Upsilon,\mu\Psi(n),\sigma^{2})\right)+n\rho\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})\right]$
$\overset{(b)}{<}$ $a(\Upsilon,\rho)e^{-b(\Upsilon,\rho)n}\,\ \ \,\text{for some}\,\ \,a(\Upsilon,\rho),b(\Upsilon,\rho)>0$ (87)
Note that the above inequality (III-B) holds for any $\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]$. Let $a=\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\max}\{a(\Upsilon,\rho)\}$ and $b=\underset{\Upsilon\in[\mu^{2}\Psi(n),\Psi(n)]}{\min}\{b(\Upsilon,\rho)\}$, and both $a$ and $b$ are positive from the fact that $a(\Upsilon,\rho)$ and $b(\Upsilon,\rho)$ are continuous functions of $\Upsilon$. From sphere symmetry, for any $\bm{x}$ in the support of $\bar{P}_{n}^{(n)}$, the following inequality holds.
$\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)<ae^{-bn},$ (88)
Consequently,
$\begin{split}\mathbb{P}(\mathcal{E}^{c})=&\mathbb{P}_{\bar{P}_{n}^{(n)}(\bm{X})W^{n}_{2}(\bm{Z}|\bm{X})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|\bm{X})}{{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}>\gamma_{1}\bigg)<ae^{-bn}.\end{split}$ (89)
Remark 8.
The strategy of the proof of Lemma 8 is similar as Lemma 4. We first consider the codewords with the same $l^{2}$ norm $\sqrt{n\Upsilon}$ (on the sphere with the same radius or with equal power constraint), separately. For codewords on each sphere, we can utilize the sphere symmetry to simplify the analysis. Again, the parameter $\mu$ makes the interval under control so that an universal threshold $\gamma_{1}$ can be obtained. The inner probability $\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)$ for each $\bm{x}$ is a tail probability and can be bounded by concentration inequality from Chernoff technique. As the bounds of the tail probability depends only on the radius of the codewords and are all exponentially decreasing with $n$, we can get the maximum of the tail probabilities for all $\bm{x}$, which is again exponential decreasing with $n$.
From the conclusion of Lemma 7 and Lemma 8, the first term in (72) has the following bound.
Lemma 9.
$\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|\bar{Q}^{(n)}_{\bm{Z},1})\right]\leq\log\big(1+\frac{2^{\gamma_{1}+1}}{MK}\big)+An\cdot ae^{-bn}.$ (90)
Proof.
The details can be found in (III-B), where the first step follows from (73). The inequality (a) follows by Lemma 3, the inequality (b) follows from the definition of $\mathcal{E}$, $\frac{1}{1-\iota^{\prime}}\leq 2$ for sufficiently large $n$ and Lemma 7, (c) holds because of Lemma 8 and $\log\bigg(1+\frac{2}{MK}\cdot e^{\frac{n}{2\mu}}\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)^{\frac{n}{2}}\bigg)\leq An$ for some constant $A>0$. ∎
$\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|\bar{Q}^{(n)}_{\bm{Z},1})\right$]
$\leq$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{X})\bar{P}_{n}^{(n)}(\bm{X})}\left(\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\right)$
$=$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}\right]\mathbb{P}(\mathcal{E})+\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}^{c}\right]\mathbb{P}(\mathcal{E}^{c})$
$=$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}\cdot\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MKQ^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}\right]\mathbb{P}(\mathcal{E})$
$+$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}\cdot\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MKQ^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}^{c}\right]\mathbb{P}(\mathcal{E}^{c})$
$\overset{(a)}{\leq}$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{1}{(1-\iota^{\prime})}\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MKQ^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}\right]\mathbb{P}(\mathcal{E})$
$+$ $\mathbb{E}_{W^{n}_{2}(\bm{Z}|\bm{x})\bar{P}_{n}^{(n)}(\bm{X})}\left[\log\left(1+\frac{1}{(1-\iota^{\prime})}\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MKQ^{(n)}_{\bm{Z},1}(\bm{Z})}\right)\bigg|\mathcal{E}^{c}\right]\mathbb{P}(\mathcal{E}^{c})$
$\overset{(b)}{\leq}$ $\log\big(1+\frac{2^{\gamma_{1}+1}}{MK}\big)\cdot\mathbb{P}(\mathcal{E})+\log\bigg(1+\frac{2}{MK}\cdot e^{\frac{n}{2\mu}}\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)^{\frac{n}{2}}\bigg)\cdot\mathbb{P}(\mathcal{E}^{c})$
$\leq$ $\log\big(1+\frac{2^{\gamma_{1}+1}}{MK}\big)+\log\bigg(1+\frac{2}{MK}\cdot e^{\frac{n}{2\mu}}\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)^{\frac{n}{2}}\bigg)\cdot\mathbb{P}(\mathcal{E}^{c})$
$\overset{(c)}{\leq}$ $\log\big(1+\frac{2^{\gamma_{1}+1}}{MK}\big)+An\cdot ae^{-bn}.$ (91)
At this end, we have the following upper bound of $\mathbb{E}\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\right$] by upper bounding the second and third terms of (72).
Theorem 5.
$\begin{split}\mathbb{E}&\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\right]\leq\log\big(1+\frac{2^{\gamma_{1}+1}}{MK}\big)+An\cdot ae^{-bn}\\ +&\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})\end{split}$ (92)
Proof.
Since we have bounded the first term in Lemma 9, we just consider the second and the third term in (72). For the second term, as the distributions $\bar{Q}^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},1}$ are independent of any codeword $\bm{x}$, we have
$\mathbb{E}$ $\left[\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}\right]$
$=$ $\int_{\bm{z}}\left(\int_{\bm{x}}\frac{1}{MK}\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{x}_{sw})d\bar{P}_{n}^{(n)}\right)$
$\times$ $\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}$
$\overset{(a)}{=}$ $\int_{\bm{z}}\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}$
$=$ $\mathbb{D}(\bar{Q}^{(n)}_{\bm{Z},1}\|{Q}^{(n)}_{\bm{Z},1})$. (93)
In step (a), we have used the fact that each $\bm{x}_{sw}$ are i.i.d generated from $\bar{P}_{n}^{(n)}$ and the fact that the induced distribution of $\bar{P}_{n}^{(n)}$ is $\bar{Q}^{(n)}_{\bm{Z},1}$. For the third term in (72), since the distributions $Q^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},0}$ are independent of $\bm{x}$, we have
$\mathbb{E}$ $\left[\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}\right]$
$=$ $\int_{\bm{z}}\left(\int_{\bm{x}}\frac{1}{MK}\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{x}_{sw})d\bar{P}_{n}^{(n)}\right)$
$\times$ $\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$
$\overset{(a)}{=}$ $\int_{\bm{z}}\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$
$=$ $\int_{\bm{z}}\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}\cdot Q^{(n)}_{\bm{Z},1}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$
$\overset{(b)}{\leq}$ $(1+\iota^{\prime})\int_{\bm{z}}Q^{(n)}_{\bm{Z},1}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$
$=$ $(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}(\bm{z})\|Q^{(n)}_{\bm{Z},0}(\bm{z})).$ (94)
In the above derivation, the step (a) follows from $\bar{Q}^{(n)}_{\bm{Z},1}$ is the induced distribution of $\bar{P}_{n}^{n}$ through the channel between Alice and Willie, step (b) follows from Lemma 3.
From the above (III-B), (III-B) and (III-B), we have
$\mathbb{E}$ $\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\right$]
$\leq$ $\log\big(1+\frac{2^{\gamma_{1}+1}}{MK}\big)+An\cdot ae^{-bn}$
$+$ $\mathbb{D}(\bar{Q}^{(n)}_{\bm{Z},1}\|{Q}^{(n)}_{\bm{Z},1})+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
$\overset{(a)}{\leq}$ $\log\big(1+\frac{2^{\gamma_{1}+1}}{MK}\big)+An\cdot ae^{-bn}$
$+$ $\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$ (95)
where step (a) follows from data processing inequality and the fact that $\bar{Q}^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},1}$ are the induced output distributions of $\bar{P}_{n}^{(n)}$ and $P^{(n)}_{n}$ through the channel $W_{2}^{n}$, respectively. ∎
Remark 9.
In our previous work [23], the encoder adopts sequential random coding (Theorem 1 in [23]) so that the codewords are not i.i.d generated, which makes it difficult to perform resolvability analysis. In this work, the codewords are i.i.d sampled from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$ that is close to Gaussian family. The resolvability analysis on bounding the divergence between the induced output distribution of the codewords and the distribution of noise depends much on this fact.
Remark 10.
In the resolvability analysis, we directly analyze the expectation of $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ by leveraging intermediate distributions $Q^{(n)}_{\bm{Z},1}$ and $\bar{Q}^{(n)}_{\bm{Z},1}$. The strategy is different from the technique path in DMCs [22], where the divergence between the induced distribution of the code family and the distribution of noise is first bounded using modified triangle inequality, and then each term of the upper bound is separately bounded. This is because that the modified triangle inequality does not hold for KL divergence for distributions in $\mathbb{R}^{n}$.
In the upper bound of (92), the first term represents effect of randomly sampling finite codewords. We can see that when there are more codewords, the ratio of it in the upper bound will be smaller. Moreover, the power level of the codewords and the noise at the adversary are also involved with the first term, we will explain it in detail later. The second term represents the effect of truncation in the generating distribution with respect to isotropic Gaussian distributions, which is the cost for necessary power constraint of finite blocklength random coding. The third term represents KL divergence between two isotropic Gaussian distributions, one of which is the distribution of noise, For the third term, the parameters $\Psi(n)$ and $\sigma$, which represent the power of the codewords and the power of the channel noise of $W_{2}$, are the key elements for the coding scheme and will be determined later.
III-C Existence of the Codebooks and Achievability Bound
In last two sections, we have established the decoding error bounds in (62), (68) and the resolvability bounds in (92). Although the bound on error probability of the first type in (62) is in the form of maximal probability of error, the second type of error probability in (68) and the resolvability bound are the average over the generating distribution $\bar{P}_{n}^{(n)}$. In this section, we will first combine these results to show that, with positive probability, a new family of codebooks can be obtained by an expurgation operation on the previous codebooks sampled from $\bar{P}_{n}^{(n)}$. Each codeword in this new family has an upper bound on the second type of decoding error $\varepsilon^{(2)}$, at the mean time, the induced distribution of the whole family is close to the distribution of the channel noise. As a result, the maximal decoding error probability of each codeword of this new family is low, while the induced distribution of the whole family satisfies the covert constraint. We determined various intermediate parameters and provide the size of the codebooks as well as the number of key bits. The roadmap is the same as [22]. Nevertheless, as we are dealing with the AWGN channel, the technical details are different. As we will see, the noise level $\sigma^{2}$ and the hyper-parameter of truncation $\mu$ play a vital role, which are absent from the case of DMCs.
We first recall McDiarmid’s inequality below.
Theorem 6.
(McDiarmid’s inequality): Let $\{X_{i}\}_{i=1}^{n}$ be independent random variables on some measurable space $\mathcal{X}$, and a measurable function $U=f(X_{1},\cdots,X_{n})$ with $f:\mathcal{X}^{n}\rightarrow\mathbb{R}$ satisfying the following bounded difference assumption:
$\underset{x_{1},\cdots,x_{n},x^{\prime}_{i}\in\mathcal{X}}{\sup}$ $|f(x_{1},\cdots,x_{i},\cdots,x_{n})$
- $f(x_{1},\cdots,x^{\prime}_{i},\cdots,x_{n})|\leq d_{i}$ (96)
for every $1\leq i\leq n$, where $d_{1},\cdots,d_{n}$ are some nonnegative real constants. Then, for every $r>0$,
$\mathbb{P}\big(U-\mathbb{E}U\geq r\big)\leq\exp\bigg(-\frac{2r^{2}}{\sum_{i=1}^{n}d_{i}^{2}}\bigg).$ (97)
The following lemma is necessary for the application of McDiarmid’s Inequality for the codewords over the AWGN channel, the proof can be found in Appendix E.
Lemma 10.
Let $\mathcal{C}_{1}\triangleq\left\{\bm{x}_{1},\cdots,\bm{x}_{k},\cdots,\bm{x}_{M}\right\}$ and $\mathcal{C}_{2}\triangleq\{\bm{x}_{1},\cdots,\bm{x}^{\prime}_{k},\cdots,\bm{x}_{M}\}$ be two codebooks of codewords with length $n$, which have only one difference on $k$-th codeword. The codewords are of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are all in the support set of $\bar{P}_{n}^{n}$. If $Q^{(n)}_{\bm{Z},\bm{C}_{1}}$ and $Q^{(n)}_{\bm{Z},\bm{C}_{2}}$ are the mixture Gaussian distributions induced by $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ through the channel $W_{2}$, respectively. The following bound holds for the difference between $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{1}}\|Q^{(n)}_{\bm{Z},0})$ and $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{2}}\|Q^{(n)}_{\bm{Z},0})$.
$\begin{split}\big|\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{1}}\|Q^{(n)}_{\bm{Z},0})&-\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{2}}\|Q^{(n)}_{\bm{Z},0})\big|\leq\frac{4n\Psi(n)}{M\sigma^{2}}.\end{split}$ (98)
Regarding the quantity $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ as a function of the codewords $\bm{x}_{sw},s=1,\cdots,K,w=1,\cdots,M$, we apply McDiarmid’s Inequality with Lemma 10 and obtain the following lemma. As the proof is simple, we omit it here.
Lemma 11.
Consider the AWGN channel $W_{2}$ defined in (6), if $\{\bm{x}_{w},w=1,\cdots,M\}$ are $M$ codewords randomly sampled from $\bar{P}_{n}^{(n)}$ and $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{Z})=\frac{1}{M}\sum_{w=1}^{M}W_{2}(\bm{Z}|\bm{x}_{w})$ is the induced mixture distribution of the codebook through the channel $W_{2}$, then for any $\lambda>0$, the following inequality holds,
$\mathbb{P}\bigg(\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})-\mathbb{E}\big(\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\big)\geq\lambda\bigg)$
$\leq\exp\bigg(-\frac{M\lambda^{2}\sigma^{4}}{8n^{2}\Psi^{2}(n)}\bigg).$ (99)
The following lemma is almost a restatement of Lemma 4 in [22]. It provide a sufficient condition that we can expurgate part of the codewords so that the remaining codewords constitute a family of codebooks such that its induced distribution is close to the distribution of noise, at the same time, the average decoding error probability of the second type for each of the codebook is low.
Lemma 12.
With AWGN channels $W_{1}$ and $W_{2}$ defined in (5) and (6), respectively, let the codewords $\bm{x}_{sw},s=1,\cdots,K,w=1,\cdots,M$ be i.i.d sampled from $\bar{P}_{n}^{(n)}$, and denote $\epsilon^{(1)}_{sw}$ and $\epsilon^{(2)}_{sw}$ as the decoding error probability of the first type and second type as (42) and (43). For any subset $\mathcal{I}\subset\langle M\rangle\times\langle K\rangle$, define a mixture distribution $Q_{\bm{Z},\bm{C}}^{\mathcal{I}}$ as
$Q_{\bm{Z},\bm{C}}^{\mathcal{I}}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(s,w)\in\mathcal{I}}W^{n}_{2}(\bm{Z}|\bm{x}_{sw})$ (100)
For this mixture distribution and every positive $\lambda_{1},\lambda_{2},\lambda_{3}$, define events
$\begin{split}\mathcal{E}_{1}\triangleq\left\{\underset{s}{\max}\frac{1}{M}\sum_{w=1}^{M}\varepsilon_{sw}^{(2)}\leq\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\right\}\end{split}$ (101)
and
$\begin{split}\mathcal{E}_{2}\triangleq&\left\{\forall\mathcal{I}\subset\langle K\rangle\times\langle M\rangle\,\ \,\text{with}\,\ \,|\mathcal{I}|=\lambda_{2}MK,\right.\\ \mathbb{D}(Q_{\bm{Z},\bm{C}}^{\mathcal{I}}\|Q^{(n)}_{\bm{Z},0})&\leq\log\big(1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}MK}\big)+An\cdot ae^{-bn}\\ +&\left.\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})+2\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})+\lambda_{3}\right\}.\end{split}$ (102)
Then $\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})>0$ if
$\begin{split}\frac{1}{\lambda_{1}}+\exp\bigg[-M\cdot\bigg(\frac{\lambda_{2}\lambda_{3}^{2}\sigma^{4}}{8n^{2}\Psi^{2}(n)}-\mathbb{H}_{b}(\lambda_{2})\bigg)\bigg]<1.\end{split}$ (103)
Proof.
From the same arguments in [22], the probability of the first event satisfy
$\mathbb{P}(\mathcal{E}_{1})$ (104)
$=$ $\mathbb{P}\bigg(\underset{s}{\max}\frac{1}{M}\sum_{w=1}^{M}\varepsilon_{sw}^{(2)}\leq\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\bigg)$
$=$ $\bigg(1-\mathbb{P}\big(\frac{1}{M}\sum_{w=1}^{M}\varepsilon_{sw}^{(2)}\geq\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\bigg)^{K}$
$\geq$ $(1-\frac{1}{\lambda_{1}})^{K}.$ (105)
For the second event, the following bound follows from McDiarmid’s inequality,
$\begin{split}\mathbb{P}(\mathcal{E}_{2})&\geq 1-\dbinom{MK}{\lambda_{2}MK}\times\exp\bigg(-\frac{\lambda_{2}\lambda_{3}^{2}MK\sigma^{4}}{8n^{2}\Psi^{2}(n)}\bigg)\\ \geq&1-\exp(MK\mathbb{H}_{b}(\lambda_{2}))\times\exp\bigg(-\frac{\lambda_{2}\lambda_{3}^{2}MK\sigma^{4}}{8n^{2}\Psi^{2}(n)}\bigg)\end{split}$ (106)
Therefore, we obtain
$\begin{split}&\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})\\ =&1-\mathbb{P}(\mathcal{E}^{c}_{1}\cup\mathcal{E}^{c}_{2})\\ \geq&1-\mathbb{P}(\mathcal{E}^{c}_{1})-\mathbb{P}(\mathcal{E}^{c}_{2})\\ =&\mathbb{P}(\mathcal{E}_{1})+\mathbb{P}(\mathcal{E}_{1})-1\\ \geq&(1-\frac{1}{\lambda_{1}})^{K}-\exp(MK\mathbb{H}_{b}(\lambda_{2}))\times\exp\bigg(-\frac{\lambda_{2}\lambda_{3}^{2}MK\sigma^{4}}{8n^{2}\Psi^{2}(n)}\bigg)\end{split}$ (107)
Hence, $\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})>0$ if
$\begin{split}1-\frac{1}{\lambda_{1}}\geq\exp\bigg[-M\cdot\bigg(\frac{\lambda_{2}\lambda_{3}^{2}\sigma^{4}}{8n^{2}\Psi^{2}(n)}-\mathbb{H}_{b}(\lambda_{2})\bigg)\bigg].\end{split}$ (108)
Remark 11.
Different from Lemma 4 in [22], here the event $\mathcal{E}_{1}$ only bounds the average error probabilities of the second type for all sub-codebooks. We have derived upper bound on the maximal decoding error probability of the first type in Lemma 4.
We have illustrated the random coding scheme, decoding method in Section III-A. The upper bound on the first type of decoding error in Lemma 4 and the average of the second type of decoding error has been analyzed in Lemma 6. We also bounded the expectation of the discriminative metric - $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$. A legitimate covert code should be based on the previous Lemma 12. The choice of various parameters both reliability and covertness will be illustrated in our main results as next theorem to realize the event $\mathcal{E}_{1}\cap\mathcal{E}_{2}$. We will provide characterizations on the size $M_{n}$ of each codebook for covert communication over AWGN channels with sufficient large $n$ through proper choice of the power level $\Psi(n)$. The theorem also characterizes the necessary number of key bits $K_{n}$, and it is shown that $K_{n}$ depends not only on the noise level $\sigma^{2}$ at the adversary’s channel and covert parameter $\delta$, but also depends on the hyper-parameter of truncation $\mu$.
Theorem 7.
For covert communication over AWGN channels $W_{1}$ and $W_{2}$ in (5) and (6), a sufficient condition for the existence of an $(M,K,n,\epsilon,\sigma,\delta)$ code is listed as the following three requirements
  1. 1.
    Reliability (maximal error probability)
    $\epsilon_{1}+\frac{\lambda_{1}}{1-\lambda_{2}}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\leq\epsilon,$ (109)
  2. 2.
    Covertness
    $\begin{split}\log&\big(1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}M_{n}K_{n}}\big)+An\cdot ae^{-bn}\\ +&\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})\\ +&\lambda_{3}\leq\delta.\end{split}$ (110)
  3. 3.
    In addition
    $\begin{split}1-\frac{1}{\lambda_{1}}\geq\exp\bigg[-M_{n}\cdot\bigg(\frac{\lambda_{2}\lambda_{3}^{2}\sigma^{4}}{8n^{2}\Psi^{2}(n)}-\mathbb{H}_{b}(\lambda_{2})\bigg)\bigg].\end{split}$ (111)
is listed as follows.
  1. (a)
    Choose intermediate parameters $\lambda_{1},\lambda_{2},\lambda_{3}$ as (1): $\lambda_{1}=n^{c_{1}}$, (2) $\lambda_{2}=1-n^{-c_{2}}$, (3): $\lambda_{3}=n^{-c_{3}}$ such that $c_{1},c_{2},c_{3}$ all positive, meanwhile, they satisfy $0<c_{1}<c_{0}$, $c_{1}+c_{2}<c_{0}$ and $0<2c_{3}<c_{2}-1$.
  2. (b)
    $\begin{split}\log M_{n}=&nC(\mu,\Psi(n))+n\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))\\ +&\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}\\ \times&\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})\\ +&O(\log n).\end{split}$ (112)
  3. (c)
    The sizes of the codebooks and the key satisfy:
    $\frac{2^{\gamma_{1}+1}}{M_{n}K_{n}}=\frac{1}{n^{c_{5}}}$ (113)
    with some $c_{5}>0$, where
    $\gamma_{1}=n\mathsf{C}(\mu,\Psi,\sigma^{2})-n(1+\rho)\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})$ (114)
    with any $\rho>0$.
  4. (d)
    The power level $\mu\Psi(n)=c_{4}\cdot\frac{\sqrt{n\delta}}{n}$ with $c_{4}$ is any constant such that
    $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})<\delta.$ (115)
Under the above conditions, with sufficiently large $n$, we can choose $c_{0}=13$, $c_{1}=1$, $c_{2}=11$, $c_{3}=4$, and $c_{5}=3$ and a proper power level $c_{4}=\frac{2\sigma^{2}}{\mu}\sqrt{\delta\ln 2}$. We obtain We obtain (1) and the number of key bits should satisfy (21).
Proof.
Based on Lemma 12, for each index $s$, we expurgate the $(1-\lambda_{2})M_{n}$ codewords with largest $\varepsilon_{sw}^{(2)}$ for from each sub-codebook $\{\bm{x}_{sw},w\in\langle M\rangle\}$. Thus,we get a new family of codebooks, each of which has size $\lambda_{2}M$. Without loss of generality, we assume that the remaining codewords have indexes from $1$ to $\lambda_{2}M$, and denote the maximum of error probability of the second type in these new codes as $\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}$. These codewords with indexes from $\lambda_{2}M+1$ to $M$ have error probabilities of the second type at least larger than $\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}$. As the average error probability of the second type for each old codebook is less than $\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$. We have for each $s$, the following inequality holds,
$\begin{split}&\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\cdot M_{n}\\ \geq&\sum_{w=1}^{M}\varepsilon_{sw}^{(2)}=\sum_{w=1}^{\lambda_{2}M}\varepsilon_{sw}^{(2)}+\sum_{w=\lambda_{2}M+1}^{M}\varepsilon_{sw}^{(2)}\\ \geq&(1-\lambda_{2})M_{n}\cdot\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}.\end{split}$ (116)
Thus, the new family of codebooks should satisfy
$\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(2)}\leq\frac{\lambda_{1}}{1-\lambda_{2}}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg).$ (117)
For the condition of reliability, for sufficiently large $n$, the following inequality on the maximal error probability of the code family holds,
$\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(2)}\leq$ $\frac{1}{n^{c_{0}-c_{1}-c_{2}}}\cdot\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$
$\overset{(a)}{\leq}$ $\frac{2}{n^{c_{0}-c_{1}-c_{2}}}$
$\leq$ $\epsilon-\epsilon_{1}$ (118)
where the step (a) follows from Lemma 3. Thus, we get the maximal probability of error satisfies
$\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(2)}+\varepsilon_{sw}^{(1)}\leq\epsilon-\epsilon_{1}+\epsilon_{1}<\epsilon.$ (119)
from the conclusion of Lemma 4. For the covertness condition in (110), from (115) and Lemma 3, the term $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$ satisfies
$(1+\iota^{\prime})$ $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
$\leq$ $\delta-[\frac{2}{n^{c_{5}}}+An\cdot ae^{-bn}+\log\frac{1}{\Delta}+\frac{1}{n^{c_{3}}}]$ (120)
for sufficiently large $n$, we obtain
$\log$ $\big(1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}MK}\big)+An\cdot ae^{-bn}$ (121)
$+$ $\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})+\lambda_{3}+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
$\leq$ $\frac{2}{n^{c_{5}}}+An\cdot ae^{-bn}+\log\frac{1}{\Delta}$
$+$ $\frac{1}{n^{c_{3}}}+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
$\leq$ $\delta.$ (122)
For the third condition, we have
$\frac{\lambda_{2}\lambda_{3}^{2}\sigma^{4}}{8n^{2}\Psi^{2}(n)}-\mathbb{H}_{b}(\lambda_{2})$
$\geq$ $\frac{\mathcal{\nu}_{1}}{n^{1+2c_{3}}}-\mathbb{H}_{b}(\lambda_{2})$
$>$ $\frac{\mathcal{\nu}_{1}}{n^{1+2c_{3}}}-\frac{\nu_{2}\log n}{n^{c_{2}}}$
$>$ $\nu_{3}$ (123)
with some positive constant $\nu_{3}$, where we have used the following facts: (1) $\mathbb{H}_{b}(x)=\mathbb{H}_{b}(1-x)$ (2) $\mathbb{H}_{b}(x)\leq x\log\frac{x}{e}$ with $0<x<\frac{1}{2}$. From (46) and (53), using $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$, we get
$nC(\mu,n)=$ $\frac{n}{2}\log(1+\mu\Psi(n))$
$=$ $\frac{n}{2}\log(1+\mu c_{4}\cdot\frac{\sqrt{n}}{n})$
$=$ $\nu_{5}\sqrt{n},$ (124)
$n\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))=$ $n\cdot\frac{\Upsilon-\mu\Psi(n)}{2(1+\mu\Psi(n))}\log e$
$=$ $\nu_{6}\sqrt{n}$ (125)
$\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))=$ $\frac{2\Upsilon+\mu^{2}\Psi^{2}}{2(1+\mu\Psi)^{2}}\log^{2}e$
$=$ $\nu_{7}\frac{1}{\sqrt{n}}$ (126)
Thus, from (112), we get $\log M_{n}\geq\nu_{8}\cdot\sqrt{n}$. for some $\nu_{8}>0$. Therefore, for sufficiently large $n$
$\exp$ $\bigg[-M\cdot\bigg(\frac{\lambda_{2}\lambda_{3}^{2}\sigma^{4}}{8n^{2}\Psi^{2}(n)}-\mathbb{H}_{b}(\lambda_{2})\bigg)\bigg]+\frac{1}{\lambda_{1}}$
$\leq$ $\exp\big(-2^{\nu_{8}\sqrt{n}}\cdot\nu_{3}\big)+\frac{1}{n^{c_{1}}}$
$\leq$ $1.$ (127)
With $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$, we get a lower bound of $\log M^{*}(n,\epsilon,\sigma,\delta)$ in (III-C).
$\log M^{*}(n,\epsilon,\delta)\geq$ $\frac{n}{2}\log(1+\mu c_{4}\frac{\sqrt{n}}{n})+\frac{(\mu^{2}-\mu)c_{4}\frac{\sqrt{n}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})}\cdot\log e+\sqrt{n\cdot\frac{2\mu^{2}c_{4}\frac{\sqrt{n}}{n}+\frac{c_{4}^{2}\mu^{2}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})^{2}}}$
$\times\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})\log e+O(\log n)$
$=$ $\frac{\mu c_{4}\sqrt{n}}{2}\log e+\frac{(\mu^{2}-\mu)c_{4}\sqrt{n}}{2}\log e+\mu\sqrt{c_{4}}n^{\frac{1}{4}}\Phi^{-1}(\epsilon)\log e+o(n^{\frac{1}{4}})$
$=$ $\frac{\mu^{2}c_{4}}{2}\sqrt{n}\log e+\mu\sqrt{c_{4}}n^{\frac{1}{4}}\Phi^{-1}(\epsilon)\log e+o(n^{\frac{1}{4}})$ (127)
The condition in (113) is equivalent to
$\log M+\log K=\gamma_{1}+c_{5}\log n+1$ (128)
Consequently, the least number of bits of $K$ should satisfy
$\log K$
$=$ $\max\{0,\gamma_{1}+c_{5}\log n+1-\log M\}$
$\overset{(a)}{=}$ $\max\{0,n[\mathsf{C}(\mu,\Psi(n),\sigma^{2})-C(\mu,\Psi(n))$
- $\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})-\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))]$
- $\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}$
$\times$ $\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})+c_{6}\log n\}$
$\overset{(b)}{=}$ $\max\{0,n[\mathsf{C}(\mu,\Psi(n),\sigma^{2})-C(\mu,\Psi(n))$
- $\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})-\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))]$
$+$ $\sqrt{n\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))}\times\Phi^{-1}(1-\epsilon_{1})+c_{6}\log n\}$ (129)
where (a) holds because Lemma 8 holds for any $\rho>0$, and (b) holds because $\Phi^{-1}(\epsilon)=-\Phi^{-1}(1-\epsilon)$ and $\Phi^{-1}$ is monotone increasing. The term (129) is analyzed as follows.
$\begin{split}&n[\mathsf{C}(\mu,\Psi(n),\sigma^{2})-C(\mu,\Psi(n))]\\ =&n[\frac{1}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})-\frac{1}{2}\log(1+\mu\Psi(n))]\\ =&\frac{n}{2}\log\big[1+\frac{\mu\Psi(n)(1-\sigma^{2})}{\sigma^{2}(1+\mu\Psi(n))}\big]\\ =&\frac{1}{2}\mu\big(\frac{1}{\sigma^{2}}-1\big)c_{4}\sqrt{n}\log e+O(1).\end{split}$ (130)
In addition,
$\begin{split}&n[\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})+\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))]\\ =&\frac{(\mu-1)\Psi(n)(1+\mu\Psi(n)+\mu\sigma^{2}+\mu^{2}\Psi(n))}{2(1+\mu\Psi(n))(\sigma^{2}+\mu\Psi(n))}\log e\end{split}$ (131)
By using $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$ in (130) and (131), the key size in (129) is rewritten as
$\max\{0,\frac{1}{2}\mu\big(\frac{1}{\sigma^{2}}-1\big)c_{4}\sqrt{n}\log e$
- $\frac{n(\mu-1)(1+\mu\sigma^{2})c_{4}\sqrt{n}}{2\sigma^{2}}\}$
$=$ $\max\{0,\frac{1-\mu^{2}\sigma^{2}}{2\sigma^{2}}c_{4}\sqrt{n}\log e+O(1)\}$ (132)
and the term $\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}=O(n^{1/4})$. Let $\phi(\mu,\sigma)=\frac{1-\mu^{2}\sigma^{2}}{2\sigma^{2}}$, we discuss it in different cases.
  • If $\sigma^{2}<\frac{1}{\mu^{2}}$, i.e., the channel between Alice and Willie is less noisy than the main channel, we have
    $\log K=\frac{1-\mu^{2}\sigma^{2}}{2\sigma^{2}}c_{4}\sqrt{n}\log e+O(n^{1/4}).$ (133)
  • If $\sigma^{2}\geq\frac{1}{\mu^{2}}$, then $\phi(\mu,\sigma^{2})\leq 0$ and the first order asymptotics of the optimal number of key bits should be $O(1)$.
In summery, the number of key bits should satisfy
$\log K=\begin{cases}\frac{1-\mu^{2}\sigma^{2}}{2\sigma^{2}}\cdot c_{4}\sqrt{n}\log e+O(n^{1/4}),\\ \,\ \ \ \ \ \ \ \ \ \ \ \,\text{if}\,\ \,\sigma^{2}<\frac{1}{\mu^{2}}\\ O(1)\,\ \ \ \ \ \,\text{otherwise}\end{cases}$ (134)
Finally, we consider the inequality (115)
$\delta$ $>\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
$=\frac{n}{2}\left[\frac{\mu\Psi(n)}{\sigma^{2}}-\ln(1+\frac{\mu\Psi(n)}{\sigma^{2}})\right]$
$\geq\frac{1}{4}n\frac{\mu^{2}\Psi^{2}(n)}{\sigma^{2}}$ (135)
and we get a proper $c_{4}$ as
$c_{4}=\frac{2\sigma^{2}}{\mu}\sqrt{\delta\ln 2}.$ (136)
Combing (III-C), (134) and (136), we finally get the size of the $M$ in (1) and the size of the key in (21). ∎
Remark 12.
If we apply random coding with Gaussian distribution as in the achievability part of Theorem 5 in [12] or Theorem 9.1.1 in [38], we can not guarantee that the distribution of the codebook is Gaussian since we can not use $n\rightarrow\infty$ in the finite blocklength regime. The advantages of truncated Gaussian family as the generating distribution of the codewords are as follows. Firstly, all codewords satisfying power constraint is ensured so that we can control the power level easily. Secondly, the truncated Gaussian family is close to Gaussian family, and the effect of truncation on the output distribution is convenient to manipulate to satisfy the covert constraint. The support set of the codewords is $\{\bm{x}:\sqrt{\mu^{2}n\Psi(n)}\leq\|\bm{x}\|\leq\sqrt{n\Psi(n)}\}$. The upper bound of $\|\bm{x}\|$ is for the power constraint, and we set the inner bound $\sqrt{\mu^{2}n\Psi(n)}$ here for upper bounding the error probability of the first type. If there exists codewords with too small $l^{2}$ norm, then there exists no $\gamma$ to upper bound the first type of error in decoding. In fact, the bounding of the first type of error in Lemma 4 relies on the fact that the power of each codeword lies in a finite interval (See Remark 5). Here we choose truncated Gaussian with decreasing variance $\mu\Psi(n)$. The parameter $\mu$ acts as both the parameter of variance and the parameter of truncation, which makes the analysis more convenient and the analytic formulas more concise. The truncation parameters for the inner bound and outer bound of $l^{2}$ norm of the codewords can be different, which will lead to the first and second order asymptotics with different coefficients. Nevertheless, it will not change the order of them as $O(n^{\frac{1}{2}})$ and $O(n^{\frac{1}{4}})$. In addition, the truncation parameter of the outer bound of radius will be reflected in the coefficient of these asymptotics as the power constraint is necessary in AWGN channels.
Remark 13.
To satisfy the covert constraint with KL divergence in (12), the power level should be $\Psi(n)=O(\frac{1}{\sqrt{n}})$ in (115) of the condition (d). The kernel lies in the divergence between two Gaussian distributions $Q^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},0}$. If the covert constraint is in the form of other divergences with proper upper bound $\delta$, such as total variational distance, Hellinger distance or $\chi^{2}$-divergence, the results will be almost the same except that $\Psi(n)$ may have a different constant coefficient. This fact will be illustrated later in Section IV-C.
Remark 14.
The differences between Theorem 7 here and Theorem 5 in Section V of [12] are as follows. Firstly, the main channel $W_{1}$ and the channel at the adversary $W_{2}$ defined in (5) and (6) here are regarded as the same in [12] (with the same level of background noise $\sigma^{2}$). Hence, the defined $L$ there is $1$, irrespectively of the power level $\sigma^{2}$. On the contrary, the first order and second order asymptotics depend on $\sigma^{2}$ in our setting because we allow the noise levels of (5) and (6) to be different. Intuitively, higher level of background noise at the adversary allows the power level of the sending signal to be higher, which leads to larger capacity. Secondly, our analysis is in the finite blocklength regime, where $n$ can only be sufficiently large and $\epsilon$ is usually a constant such as $10^{-3}$. Thus, we can not utilize $n\rightarrow\infty$ or $\epsilon\rightarrow 0$ as the definition of $L$ in (7) in [12].
Remark 15.
Our result on the size of key bits is slightly different from the case in discrete memoryless channels [22], and is also different from Theorem 6 in [10] where the numbers of key bits are $O(1)$ if the adversary is less noisy ($D_{Q}\leq D_{P}$). In this setting, the number of key bits also depends on the parameter of truncation $\mu$ with finite blocklength over AWGN channels. Even if the channel between Alice and Willie is severer than the main channel, the number of key bits should be of $O(\sqrt{n})$ if the noise level is only a bit larger. This is the necessary cost of power constraint for the sending codewords with finite blocklength over AWGN channels, which will lead to the truncation for Gaussian random coding in the finite blocklength regime. The truncation is not necessary when we only consider the first order asymptotics, such as the proof of the achievability part of Theorem 5 in [12]. This can be reflected by the following fact. If we let $\mu=\mu(n)=1-\frac{1}{n^{\tau}}$ with $0<\tau<\frac{1}{2}$, the behavior of $\Delta$ can be found as following Fig. 4. In this case, Corollary 1 still holds for sufficiently large $n$ and the other lemmas and theorems also hold. Under the circumstance, the codewords are far away from with equal power when $n$ is finite. The truncation effect $\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})$ on the output distribution can be neglected in the expectation of the divergence $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ when $n$ is sufficiently large. Consequently, the first order and the second order asymptotics of the throughput are $\sigma^{2}(\delta n\log e)^{1/2}$ and $\sigma^{2}(\delta n\log e)^{1/2-\tau}$, respectively. In addition, the first- and second order asympotics of the number of key bits are $(1-\sigma^{2})\sqrt{n\delta\log e}$ and $O(n^{\frac{1}{2}-\tau})$, respectively. If we further let $\tau=\frac{1}{2}$, we can see that $\Delta$ is stationary with respect to $n$ from Fig .4 and can not be neglected. From the definition of $\bar{g}^{(n)}(\bm{x})$ and $g^{(n)}(\bm{x})$ and their relationship with $\bar{g}_{1}^{(n)}(\bm{y})$$g_{1}^{(n)}(\bm{y})$ and $\bar{h}_{1}^{(n)}(\bm{z})$$h^{(n)}_{1}(\bm{z})$, we can get that $\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}$ and $\frac{\bar{h}_{1}^{(n)}(\bm{z})}{h^{(n)}_{1}(\bm{z})}$ are bounded for all $n$. From (68), the average of the error probability of the second type can be sufficiently small with same $M_{n}$. Meanwhile, Lemma 4 for the maximum of error probability of the first type still holds. Nevertheless, from the covertness condition (110), the covert condition holds only for $\{\delta:\delta>\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})=-\log\Delta\}$ since we can choose $\Psi(n)$ such that the other terms in (110) are sufficiently small for sufficiently large $n$. For these $\delta$, the second order asymptotics of $\log M^{*}(n,\epsilon,\sigma,\delta)$ is $O(n^{\frac{1}{4}})$. The additional cost on the covertness stems from the truncation. The above analysis implies that it is preferable to choose $\mu$ close to $1$ when the blocklength $n$ is large and a moderate $\mu$ between $\frac{1}{2}$ and $1$ when we have a moderate blocklength $n$. Nevertheless, as the blocklength $n$ is always finite, $\mu<1$ is necessary, which implies that the transmitter still needs keys to assist when $1<\sigma^{2}\leq\frac{1}{\mu^{2}}$.
Fig. 4: The value of $\Delta$ with proper chosen $\mu(n)=1-\frac{1}{n^{\tau}}$, which is from numerical computation of the analytic formula (32).
Our results reveal that the consumption of covertness comes from two aspects. The first is average power of the signal. From Theorem 5 of [12], Gaussian signals will result in least consumption of covertness. Nevertheless, we can not obtain perfect Gaussian signals in the finite blocklength regime. There is always some gap between the empirical distribution of signals and the Gaussian distribution. The second ingredient is that practical signals have finite power, which is the reason for truncation. These two aspects are reflected in two KL divergence terms in the inequality (110) as $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$ and $\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})$, respectively. The key parameter of the two terms are $\mu$ and $\Psi(n)$. We will discuss on the choice of $\Psi(n)$ in Section IV later.
We can see that both of the sizes of sub-code and the key depend on $\mu$, $\sigma$ and $\delta$. The first order of $\sigma$ and $\delta$. The first order of $\log M^{*}(n,\epsilon,\sigma,\delta)$ increases with $\mu$, $\sigma^{2}$ and $\delta$, which is consistent with the intuition. With fixed $\mu$ and $\delta$, when the noise level at the channel between Alice and Willie is high, then Alice can increase the power of the code and thus the size of the code without violate the covert constraint and maintain the same reliability. The same reasoning applies for fixed $\mu$ and $\sigma$. For the size of key, we can see that with fixed $\delta$, a larger $\sigma$ will lead to a smaller size of key. This is because the transmitter needs less sub-codes to confuse the adversary if the adversary has higher level of background noise.
III-D Converse Bound
In this section, the converse bound under covert constraint (12) is investigated, and the technique is to reconcile the bound in terms of the output distribution of the channel to get an upper bound of average power level of the codebook, which has been utilized in [25][12]. The difference is that the average power considered here should be over all messages and keys for $n$ channel uses. To prove the converse for any code corresponding to any key, we follow the argument in Theorem 7 of [21] and leverage the pigeon-hole principle in our arguments.
First, following the inequality (231) in [21], we have
$\delta\geq\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\geq n\mathbb{D}(\bar{Q}^{(n)}_{\bm{Z},\bm{C}}\|Q_{\bm{Z},0}),$ (137)
where $\bar{Q}^{(n)}_{\bm{Z},\bm{C}}=\frac{1}{n}\sum_{i=1}^{n}\sum_{s=1}^{K}\sum_{w=1}^{M}\frac{1}{MK}W_{2}(\bm{z}(i)|\bm{x}_{sw}(i))$ is the average output distribution at the adversary per channel use over all the messages and keys and $Q_{\bm{Z},0}$ is the distribution of noise at the adversary. Following the same arguments around (74) in [12], we have the following lemma.
Lemma 13.
For any $(M,K,n,\epsilon,\Psi)_{a}$ code which satisfies the covert constraint, then the corresponding average power $\Psi$ should be less than or equal to the solution $x$ of the equation
$\mathbb{D}(\bar{\mathbb{Q}}_{G}\|Q_{\bm{Z},0}))=\frac{\delta}{n},$ (138)
where $\bar{\mathbb{Q}}^{(n)}_{G}=\mathcal{N}(\bm{0},(\sigma^{2}+\varpi)\bm{I}_{n}).$
Denote the solution of (138) as $\Psi(\delta,\sigma,n)$. From Lemma 13, it is an upper bound of the average power of any $(M,K,n,\epsilon,\sigma,\delta)$ code . Note that the average power considered here is over all messages and keys with $n$ channel uses. To prove the converse bound for any code, we still need the following corollary.
Corollary 3.
There exists a subset $\mathcal{D}$ of codewords $\bm{x}_{sw},s=1,\cdots,K,w=1,\cdots,M$ of size $MK/n$ such that
$\underset{\|\bm{x}_{sw}\|^{2}\leq n\Psi(n)}{\max}\Psi(n)\leq\frac{n}{n-1}\Psi(\delta,\sigma,n)$ (139)
Proof.
If we sort the power of the codewords from maximum to minimum, let $\mathcal{D}$ be the last $\frac{MK}{n}$ codewords. We claim that these codewords have maximal power less than or equal to $\frac{n}{n-1}\Psi(\delta,\sigma,n)$. Otherwise, all the remaining codewords have power at least $\frac{n}{n-1}\Psi(\delta,\sigma,n)$, then the average power of the whole $MK$ codewords is larger than
$\frac{1}{MK}\underset{\bm{x}\in\mathcal{D}^{c}}{\sum}\frac{n}{n-1}\Psi(\delta,\sigma,n)$
$=$ $\frac{1}{MK}\cdot\frac{(n-1)MK}{n}\cdot\frac{n}{n-1}\Psi(\delta,\sigma,n)$
$=$ $\Psi(\delta,\sigma,n),$ (140)
which is a contradiction. ∎
The above corollary show that there exists “part” of the whole codewords such that the maximal power is bounded. By using existing converse bounds under maximal power constraint, we can upper bound the size of that part, and hence the size of each subcode, which lead to the converse part of Theorem 1.
Theorem 8.
For covert communication over AWGN channels (5) and (6), any $(M,K,n,\epsilon,\sigma,\delta)$ code should satisfy (1).
Proof.
For any $(M,K,n,\epsilon,\sigma,\delta)$ code, we proceed with the arguments in Corollary 13 and Corollary 3. From the pigeon-hole principle, there exists at least one sub-code $\mathcal{C}^{s}$ such that
$|\mathcal{D}\cap\mathcal{C}^{s}|\geq M/n$ (141)
and all the codewords in $\mathcal{D}\cap\mathcal{C}^{s}$ have maximal power less than $\frac{n}{n-1}\Psi(\delta,\sigma,n)$. Thus, the size of $\mathcal{D}\cap\mathcal{C}^{s}$ must satisfy
$\log|\mathcal{D}\cap\mathcal{C}^{s}|$ $\leq\log M^{*}_{m}(n,\epsilon,\frac{n}{n-1}\Psi(\delta,\sigma,n)),$ (142)
Using a converse bound under maximal power constraint with varying power ( the formula (24) in [23]), we get
$\log|\mathcal{D}\cap\mathcal{C}^{s}|\leq$ $n\log(1+\frac{n}{n-1}\Psi(\delta,\sigma,n))$
- $\sqrt{n\mathsf{V}(\frac{n}{n-1}\Psi(\delta,\sigma,n))}$
$\times$ $\Phi^{-1}(1-\epsilon_{1})+O(\log n),$ (143)
where we have used $\Phi^{-1}(1-\epsilon_{1})$ to substitute [^2] Here $Q(x)=\int_{x}^{\infty}\frac{1}{\sqrt{2\pi}}e^{-t^{2}/2dt}$ $Q^{-1}(\epsilon)$. From the fact that for any $\eta>1$
$\frac{1}{2}[x-\ln(1+x)]\geq\frac{x^{2}}{4\eta},$ (144)
we get an upper bound on $\Psi(\delta,n)$ as $\Psi_{u}(\delta,n)=\sqrt{\frac{4\eta\delta\ln 2}{n}}\cdot\sigma^{2}$ and from (141) and (III-D) we further have
$\begin{split}\log&M^{*}(n,\epsilon,\sigma,\delta)\\ \leq&\frac{n}{2}\log\left(1+\frac{n}{n-1}\cdot\sqrt{\frac{4\eta\delta\ln 2}{n}}\cdot\sigma^{2}\right)\\ +&\sqrt{\frac{n\log^{2}e}{2}\cdot V_{2}(n)}\Phi^{-1}(\epsilon)+O(\log n),\end{split}$ (145)
with
$V_{2}(n)=\frac{\sigma^{4}\delta\eta\ln 2+\sqrt{\sigma^{4}\delta\eta n\ln 2}}{\sigma^{4}\delta\eta\ln 2+\sqrt{\sigma^{4}\delta\eta n\ln 2}+\frac{n}{4}},$ (146)
[…]
Remark 16.
The argument of the converse bound is similar as [22], where an upper bound on the weight of the codewords is related to further establish the converse bound for covert binary-input discrete memoryless channels. In this paper, we choose KL divergence as the covertness metric because it is relative easy to establish the optimality of Gaussian distributions in the converse part. The optimality cannot be granted when the covert metric is total variational distance or other divergences.
IV On the Choice of Truncated Gaussian Distribution in Random Coding
In Section III-A, truncated Gaussian distribution is adopted as the generating distribution of the AWGN covert codes for achievability in Theorem 1. From the comparison between the achievability and the converse in Theorem 1, we can see that the generating distribution is optimal in the sense that it leads to nearly maximal coding rate when $n$ is large. In this section, the reason of choosing truncation Gaussian for random coding is investigated. We first explore it from a viewpoint of local information geometry, then we further analyze it by considering the asymptotic behavior of several divergences. The viewpoints in the section helps us to understand the underlying principle of choosing generating distribution for random coding in covert communication.
IV-A Quasi-$\varepsilon$-neighborhood Distributions
This section will first introduce the notion of quasi-$\varepsilon$ neighborhood. It has been used in [40][41] for local approximation of divergence in Euclidean information theory. Then the characterization of KL divergence and total variation distance for the distributions in a quasi-$\varepsilon$ neighborhood is provided. The analysis will be the base of subsequent subsections.
Definition 6.
For a given $\varepsilon>0$, the quasi-$\varepsilon$-neighborhood [41] of a reference distribution $P_{0}(z)$ on $\mathcal{Z}$ is a set of distributions in a $\chi^{2}$-divergence ball of $\varepsilon^{2}$ about $P_{0}(\bm{x})$, i.e.,
$\mathcal{N}_{\varepsilon}(P_{0})\triangleq\{P_{1}:\mathbb{\chi}^{2}(P_{1}\|P_{0})\leq\varepsilon^{2}\},$ (147)
From the above definition, for any distribution $P_{1}$ with pdf $p_{1}(\bm{x})$ in the quasi-$\varepsilon$-neighborhood of $P_{0}\in\mathbb{R}^{n}$ with pdf $p_{0}(\bm{x})$, if we define
$h_{\varepsilon}(\bm{x})=\frac{p_{1}(\bm{x})-p_{0}(\bm{x})}{\varepsilon\cdot p_{0}(\bm{x})},$ (148)
then the following lemma holds.
Lemma 14.
Any distribution $P_{1}$ in the quasi-$\varepsilon$-neighborhood of $P_{0}$ in $\mathbb{R}^{n}$ can be represented by the function $h(\bm{x})$ which satisfies $\langle p_{0},h_{\varepsilon}\rangle\triangleq\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\cdot h_{\varepsilon}(\bm{x})d(\bm{x})=0$.
Proof.
From the fact that $1=\int_{\mathbb{R}^{n}}p_{1}(\bm{x})d(\bm{x})=\int_{\mathbb{R}^{n}}p_{0}(\bm{x})d(\bm{x})+\varepsilon\cdot\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\cdot h_{\varepsilon}(\bm{x})d(\bm{x})=1+\varepsilon\cdot\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\cdot h_{\varepsilon}(\bm{x})d(\bm{x})$, the conclusion is obvious. ∎
In the following, we consider the distribution $P_{1}$ in the quasi-$\varepsilon$-neighborhood of $P_{0}$ such that the corresponding $h_{\varepsilon}(\bm{x})$ satisfies $\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h_{\varepsilon}(\bm{x})|\leq 1$. For this type of $P_{1}$ and $P_{0}$, KL divergence and TVD between them satisfy the following bounds.
Theorem 9.
For any sufficiently small $\varepsilon$, any distribution $P_{1}(\bm{x})$ in $\epsilon$-neighborhood of $P_{0}(\bm{x})$ with $h(\bm{x})$ satisfying $\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h_{\varepsilon}(\bm{x})|\leq 1$, assume that $\int_{\mathbb{R}^{n}}p_{0}(\bm{x})h^{n}_{\varepsilon}(\bm{x})d\bm{x}<\infty$ for any $n\geq 2$, we have
$V_{T}(P_{1},P_{0})\leq\frac{1}{2}\varepsilon,\,\,D(P_{1}\|P_{0})\leq\varepsilon^{2}\log e.$ (149)
The proof can be found in Appendix F.
IV-B Realization of Quasi-$\varepsilon$-neighborhood Distributions
In this section, we first introduce a method to realize quasi-$\varepsilon$-neighborhood distributions - using probability measure close to Dirac measure as convolution operator. Then, we show this method is strongly related to the optimal channel coding in covert communication over AWGN channels.
The following technical lemma is trivial and the proof is not included here.
Lemma 15.
The characteristic function for $\delta_{0}^{n}(\bm{x})$ is $1$ in $\mathbb{R}^{n}$.
In the following, we reconsider the distribution family $P_{n}^{(n)}\sim\mathcal{N}(\bm{0},\mu\Psi(n)\bm{I}_{n})$ in Section II-E and use $\|\cdot\|_{w}^{*}$ for the metric of weak convergence  [^3] Weak convergence with respect to bounded continuous functions. It has metric such as Lévy-Prokhorov metric[42].
Theorem 10.
The sequence of Gaussian distributions $P_{n}^{(n)}$ with $\Psi(n)=O(\frac{1}{n^{\tau}})$ satisfy
$\|P_{n}^{(n)}-\delta_{0}^{n}\|_{w}^{*}\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty.$ (150)
The proof can be found in Appendix G.
Remark 17.
In the above theorem, we did not prove that the sequence of Gaussian distributions converge to a single Dirac measure. In fact, with fixed $l$, Gaussian distributions $P_{l}^{(m)}\sim\mathcal{N}(\bm{0},\Psi(m)\bm{I}_{l})$ with $\Psi(m)=O(\frac{1}{m^{\tau}})$ converge to Dirac measure $\delta_{0}^{l}(\bm{x})$ in $\mathbb{R}^{l}$ as $m\rightarrow\infty$. Nevertheless, we proved that the distance of $P_{n}^{(n)}$ and $\delta_{0}^{n}$ with increasing $n$ tends to zero and the distance is in the weak sense.
Theorem 11.
The truncated Gaussian distributions $\bar{P}_{n}^{(n)}$ with $\Psi(n)=O(\frac{1}{n^{\tau}})$ satisfy:
$\|\bar{P}_{n}^{(n)}-\delta_{0}^{n}\|_{w}^{*}\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty.$ (151)
Proof.
As $\|\cdot\|_{w}^{*}$ is a metric,
$\|\bar{P}_{n}^{(n)}-\delta_{0}^{n}\|_{w}^{*}\leq\|P_{n}^{(n)}-\delta_{0}^{n}\|_{w}^{*}+\|P_{n}^{(n)}-\bar{P}_{n}^{(n)}\|_{w}^{*}.$
From Corollary 2 and Theorem 10, the conclusion is obvious. ∎
The convergence behavior of the sequence of truncated Gaussian distributions is illustrated in the distribution array in Fig. 5. Theorem 11 is demonstrated by the diagonal behavior of the distribution array. Now consider the distribution $\hat{Q}^{(n)}_{0}$ as the reference distribution $P_{0}$ in Definition 6. The realization of the quasi-$\varepsilon$-neighborhood distribution of $\hat{Q}^{(n)}_{0}$ is closely related to the approximations of Dirac measure $\delta_{0}(\bm{x})$ as operators on functions in Definition 5.
Lemma 16.
For any $\varepsilon>0$, there exists some $N(\varepsilon)$ and for any $n>N(\varepsilon)$, there exists some $h(\bm{y})$ with $\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}\,\,|h(\bm{y})|\leq 1$ such that
$\frac{\hat{f}_{1}^{(n)}(\bm{y})-\hat{f}_{0}^{(n)}(\bm{y})}{\hat{f}_{0}^{(n)}(\bm{y})}=\varepsilon h(\bm{y}).$ (152)
Proof.
Consider the distributions $\bar{P}_{n}^{(n)}$ and $\delta_{0}^{n}$ as convolution operators $\mathfrak{F}_{n}^{(n)}$ and $\mathfrak{F}_{0}^{(n)}$ on $C(\mathbb{R}^{n})$, respectively. From the property of Dirac delta function, we have $\delta_{0}^{n}*\hat{f}_{0}^{(n)}=\hat{f}_{0}^{(n)}$ and the density function $\hat{f}_{0}^{(n)}$ of $\hat{Q}^{(n)}_{0}$ belongs to $C(\mathbb{R}^{n})$. From Theorem 3 we further have
$\begin{split}&\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}|\hat{f}_{1}^{(n)}(\bm{y})-\hat{f}_{0}^{(n)}(\bm{y})|\\ =&\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}|(\mathfrak{F}_{n}^{(n)}\hat{f}_{0}^{(n)})(\bm{y})-(\mathfrak{F}_{0}^{(n)}\hat{f}_{0}^{(n)})(\bm{y})|\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty.\end{split}$
As $\log x$ is continuous, we have
$\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}|\log\hat{f}_{1}^{(n)}(\bm{y})-\log\hat{f}_{0}^{(n)}(\bm{y})|\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty.$ (153)
This is equivalent to the claim that
$\frac{\hat{f}_{1}^{(n)}(\bm{y})}{\hat{f}_{0}^{(n)}(\bm{y})}-1=\varepsilon h(\bm{y})$ (154)
holds for sufficiently large $n$ by choosing some $h(\bm{y})$ with $\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}\,\,|h(\bm{y})|\leq 1$. ∎
Lemma 16 is the consequence of Theorem 11 and Theorem 3, and it provides us a sufficient condition that $\hat{Q}^{(n)}_{1}$ is in the $\varepsilon$- neighborhood of $\hat{Q}^{(n)}_{0}$.
Fig. 5: The metric of weak convergence between $\bar{P}_{n}^{(n)}$ and $\delta_{0}^{n}$. On one hand, the distributions $\bar{P}_{l}^{(m)}$ in the $l$-th row converge to the unique Dirac measure $\delta_{0}^{l}$ as $m\rightarrow\infty$ due to the decreasing variance $P(m)$; on the other hand, the distances between $\bar{P}_{n}^{(n)}$ and $\delta_{0}^{n}$ converge to 0 as $n\rightarrow\infty$.
Corollary 4.
For any $\varepsilon>0$, there exists some $N(\varepsilon)$ and for any $n>N(\varepsilon)$, the output distribution $\hat{Q}^{(n)}_{1}$ corresponding to the input distribution $\bar{P}_{n}^{(n)}$ with $\Psi(n)=O(\frac{1}{n^{\tau}})$ is in the quasi-$\varepsilon$ neighborhood of the noise distribution $\hat{Q}^{(n)}_{0}$.
Proof.
The equation (154) implies that
$\frac{(\hat{f}_{1}^{(n)}(\bm{y})-\hat{f}_{0}^{(n)}(\bm{y}))^{2}}{\hat{f}_{0}^{(n)}(\bm{y})}=\varepsilon^{2}h^{2}(\bm{y})\hat{f}_{0}^{(n)}(\bm{y}).$ (155)
Integrate (155) in $\mathbb{R}^{n}$ will lead to
$\begin{split}&\mathbb{\chi}^{2}(\hat{Q}^{(n)}_{1}\|\hat{Q}^{(n)}_{0})\\ =&\varepsilon^{2}\int_{\mathbb{R}^{n}}h^{2}(\bm{y})\hat{f}_{0}^{(n)}(\bm{y})d\bm{y}\\ \leq&\varepsilon^{2}\cdot\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}\,\,|h^{2}(\bm{y})|\int_{\mathbb{R}^{n}}\hat{f}_{0}^{(n)}(\bm{y})d\bm{y}\\ \leq&\varepsilon^{2}.\end{split}$ (156)
Therefore, from Definition 6, we get the conclusion. ∎
From the above analysis, when the reference distribution is the distribution of noise, the truncated Gaussian distributions with decreasing variance are proper convolution operators that will lead to output distributions close to the distribution of noise. Hence, they are legitimate choice for the generating distributions of codebooks for AWGN covert channels. The following section will explain the optimality of $\tau=\frac{1}{2}$ by investigating the asymptotic behavior of some divergences between Gaussian distributions.
IV-C Square Root Law and Approximation of the Distribution of Noise
In Section III, we have witnessed that KL divergence between two Gaussian distributions with zero mean and i.i.d coordinates are the key ingredients in the evaluation of both the achievability and converse bounds of the throughput. In this section, the behavior of several divergences between two Gaussian distributions $\mathbb{P}^{(n)}_{1}\sim\mathcal{N}(\bm{0},(1+\theta_{n})\bm{I}_{n})$ and $\mathbb{P}^{(n)}_{0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$ is investigated in Subsection A, where $\theta_{n}=c\cdot n^{-\tau}$ representing the decreasing power level. This provides more evidence of the connection between Square Root Law and the convergence behaviors of several divergences when $\tau\geq\frac{1}{2}$ and $\tau<\frac{1}{2}$, respectively. Our analysis provides an explanation why we choose $\Psi(n)=O(\frac{1}{n^{\tau}})$ with $\tau=\frac{1}{2}$ in Gaussian random coding in the achievability.
IV-C1 Asymptotic behaviors of $\mathbb{V}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ and $\mathbb{D}(\mathbb{P}^{(n)}_{1}\|\mathbb{P}^{(n)}_{0})$
We will use Hellinger distance as bounds of $\mathbb{V}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ [43]:
$H^{2}(P,Q)\leq V_{T}(P,Q)\leq\sqrt{1-(1-H^{2}(P,Q))^{2}}.$ (157)
The following lemma is from the definition of the square of Hellinger distance (3).
Lemma 17.
If $\theta_{n}\sim c\cdot n^{-\tau}$ with $0<\tau<1$ where $\sim$ represents the same asymptotic order, then
$H^{2}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})\rightarrow\begin{cases}0&\tau>\frac{1}{2},\\ 1&0<\tau<\frac{1}{2},\end{cases}\,\ \,n\rightarrow\infty.$ (158)
The proof can be found in Appendix H.
Theorem 12.
If $\theta_{n}\sim c\cdot n^{-\tau}$ with $0<\tau<1$, then
$V_{T}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})\rightarrow\begin{cases}0&\tau>\frac{1}{2},\\ 1&0<\tau<\frac{1}{2},\end{cases}\,\ \,n\rightarrow\infty,$ (159)
$D(\mathbb{P}^{(n)}_{1}\|\mathbb{P}^{(n)}_{0})\rightarrow\begin{cases}0&\tau>\frac{1}{2},\\ \infty&0<\tau<\frac{1}{2},\end{cases}\,\ \,n\rightarrow\infty.$ (160)
Proof.
From the inequality (157), the asymptotic behavior of $V_{T}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ in (159) is the direct consequence of Lemma 17.
For the quantity $D(\mathbb{P}^{(n)}_{1}\|\mathbb{P}^{(n)}_{0})$, we have
$\begin{split}&D(\mathbb{P}^{(n)}_{1}\|\mathbb{P}^{(n)}_{0})\\ =&\frac{n}{2}\left[\theta_{n}-\ln(1+\theta_{n})\right]\log e\\ \sim&\frac{n\theta_{n}^{2}}{2}\log e\\ \sim&\frac{c\cdot n^{1-2\tau}}{2}\end{split}$ (161)
Hence, the asymptotic behavior of $D(\mathbb{P}^{(n)}_{1}\|\mathbb{P}^{(n)}_{0})$ in (160) is obvious. ∎
IV-C2 Square Root Law and Approximation of the Distribution of Noise
The asymptotic behavior of divergences between $\mathbb{P}^{(n)}_{1}$ and $\mathbb{P}^{(n)}_{0}$ is directly related with Square Root Law over AWGN channels [11], we will explain it from the perspective of channel resolvability as follows.
Channel resolvability problem aims at approximating a desired output distribution by using a uniform input distribution with smaller support [44][45]. In covert communication over AWGN channels, the aim is to approximate the distribution of Gaussian white noise and meanwhile maximizing the information transmitted at the main channel. The covert metric acts as the criteria of the approximation. It is well known that the mutual information at the main channel is maximized by Gaussian distribution in the asymptotic sense. Moreover, identically independent distributed Gaussian inputs will lead to Gaussian output distribution with i.i.d components, which belongs the same type of distribution. Thus, intuitively, the generating distribution of the random codes should be Gaussian. The remaining question is how to choose the variance of Gaussian distribution to satisfy the approximation criteria or covert constraint, which is given as an upper $\delta$ of some divergence. More specific, we consider approximate the output distribution $\mathbb{P}^{(n)}_{0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$ using Gaussian random codes, which are generated by the distribution $\mathcal{N}(\bm{0},\theta_{n}\bm{I}_{n})$. The parameter $\theta_{n}=c\cdot n^{-\tau}$ represents the decreasing power level. Our results in last subsection reveal the following facts.
  • If $\tau<1/2$, increasing $n$ will lead to the deviation of the induced output distribution and the distribution $\mathbb{P}^{(n)}_{1}$ and thus violate the approximation criteria.
  • With any proper $\delta$ and $\tau>\frac{1}{2}$, as $n$ increases it will definitely satisfy the requirement on the upper-bound $\delta$ imposed on the divergences between $\mathbb{P}^{(n)}_{0}$ and $\mathbb{P}^{(n)}_{1}$. However, it leaves some room for increasing $\theta_{n}$ to increase the mutual information at the main channel.
  • With given $0<\delta<1$ and $p_{n}=c\cdot n^{-\frac{1}{2}}$ where $c$ is a constant, we can increase $n$ to satisfy any small decoding error probability $\epsilon$ without violating any proper covert constraint since the divergences between $\mathbb{P}^{(n)}_{0}$ and $\mathbb{P}^{(n)}_{1}$ will be stationary and the value depends on $c$.
In this work and our previous work [23], truncated Gaussian distributions are chosen as the distributions for random coding over AWGN channels. There are two reasons of the choice. First, the way of truncation is at our choice, which will be convenient to satisfy the power constraint. The corresponding output distributions are close to Gaussian distributions with zero mean and i.i.d coordinates, hence they will lead to relatively tight normal approximations of finite blocklength throughput, which has been explained in [23]. Second, as we have explained in Section IV, truncated Gaussian distributions with proper chosen variances are good approximations of Dirac measures, which make them natural candidates on random coding for covert communications over AWGN channels. From the framework, other measures as approximations of Dirac measures can also be utilized in random coding over AWGN covert channels. Nevertheless, the number of the reliable bits from the corresponding random codes may be limited. Considering the mutual information as an approximate metric of the upper bound on the number of the transmitted bits, it is reasonable that the input distribution should be close to Gaussian.
V Conclusion
This paper presents a comprehensive analysis of covert communication over the AWGN channel, focusing on both achievability and converse bounds under finite blocklength constraints. By characterizing the first and second order asymptotics of Gaussian random coding, we establish fundamental limits on the trade-offs between reliability, covertness, and key size. Notably, the dependence of these asymptotics on the hyper-parameter $\mu$ underscores the critical role of power constraints in practical coding designs for finite $n$. Our results reveal that the required number of key bits scales as $O(\sqrt{n})$, a conclusion that remains robust even when the noise level at the legitimate receiver is marginally lower than that at the adversary. This highlights the inherent necessity of shared secrecy resources to sustain covertness in scenarios with minimal channel asymmetry. From an information-geometric perspective, the truncated Gaussian distribution emerges as the natural choice for codebook generation, balancing the trade-offs between power constraints and divergence-based covertness metrics. The Square Root Law for optimal power allocation is rigorously justified through the asymptotic behavior of divergences between Gaussian distributions and the statistical indistinguishability of noise. These findings not only unify existing theoretical frameworks but also provide practical insights for designing covert communication systems with provable guarantees.
Appendix A Proof of Lemma 1
Proof.
From Definition 3, an $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$ is necessarily an $(M,n,\epsilon,\Psi)_{m}$ code, hence the left-hand of (16) holds. An $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$ is necessarily an $(M,K,n,\epsilon,\sigma,\delta)$ code, and an $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{a}$ is necessarily a $(M,K,n,\epsilon,\sigma,\delta)$ code, hence, we have the inequalities (17) and the right-hand of (18). The inequalities on right-hand of (16) and the left-hand of (18) follow from the fact that a code which satisfies maximal power constraint less or equal than $\Psi$ necessarily has average power less or equal than $\Psi$. ∎
Appendix B Proof of Lemma 2
Proof.
Let $\Omega=\mathfrak{B}^{n}_{0}\big(\sqrt{n\Psi(n)}\big)\backslash\mathfrak{B}^{n}_{0}\big(\sqrt{\mu^{2}n\Psi(n)}\big)$, from the definition of $\bar{P}_{n}^{n}$ and $P_{n}^{n}$, we have $\frac{\bar{g}^{(n)}(\bm{x})}{g^{(n)}(\bm{x})}=\frac{1}{\Delta}$ for $\bm{x}\in\Omega$, and $\bar{g}^{(n)}(\bm{x})=0$ for $\bm{x}\in\Omega^{c}$. Hence, we have
$\begin{split}\mathbb{D}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=&\int_{\bm{x}\in\mathbb{R}^{n}}\bar{g}^{(n)}(\bm{x})\log\frac{\bar{g}^{(n)}(\bm{x})}{g^{(n)}(\bm{x})}d\bm{x}\\ =&\int_{\bm{x}\in\Omega}\bar{g}^{(n)}(\bm{x})\log\frac{1}{\Delta}d\bm{x}\\ =&-\log\Delta.\end{split}$ (162)
and
$\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=$ $\frac{1}{2}\underset{\mathbb{R}^{n}}{\int}|\bar{g}^{(n)}(\bm{x})-g^{(n)}(\bm{x})|d\bm{x}$
$=$ $\frac{1}{2}\big[\int_{\bm{x}\in\Omega}(\frac{1}{\Delta}-1)g^{(n)}(\bm{x})d\bm{x}$
$+$ $\int_{\bm{x}\in\Omega^{c}}g^{(n)}(\bm{x})d\bm{x}\big]$
$=$ $\frac{1}{2}\big[(\frac{1}{\Delta}-1)\Delta+1-\Delta\big]$
$=$ $1-\Delta.$ (163)
Appendix C Proof of Lemma 3
Proof.
Consider the distributions $\bar{P}_{n}^{(n)}$ and $P_{n}^{(n)}$ as convolution operators $\bar{\mathfrak{F}}_{n}^{(n)}$ and $\mathfrak{F}_{n}^{(n)}$ on $C(\mathbb{R}^{n})$, respectively. We apply a multivariate version of Theorem 3 for the probability density functions $f_{0}^{(n)}(\bm{y})$ and $h_{0}^{(n)}(\bm{y})$, which are in $C_{0}(\mathbb{R}^{n})$,
$\mathfrak{F}_{n}^{(n)}f_{0}^{(n)}(\bm{y})={f}_{1}^{(n)}(\bm{y}),$ (164)
$\bar{\mathfrak{F}}_{n}^{(n)}f_{0}^{(n)}(\bm{y})=\bar{f}_{1}^{(n)}(\bm{y}),$ (165)
and
$\mathfrak{F}_{n}^{(n)}h_{0}^{(n)}(\bm{z})=h^{(n)}_{1}(\bm{z}),$ (166)
$\bar{\mathfrak{F}}_{n}^{(n)}\bar{h}_{0}^{(n)}(\bm{z})=\bar{h}_{1}^{(n)}(\bm{z}).$ (167)
Corollary 2 and Theorem 3 imply the following statement.
$\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}|\bar{f}_{1}^{(n)}(\bm{y})-{f}_{1}^{(n)}(\bm{y})|\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty,$ (168)
$\underset{\bm{z}\in\mathbb{R}^{n}}{\sup}|\bar{h}_{1}^{(n)}(\bm{z})-h^{(n)}_{1}(\bm{z})|\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty.$ (169)
As the function $\varphi(x)=\log x$ is continuous, we have
$\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}|\log\bar{f}_{1}^{(n)}(\bm{y})-\log f_{1}^{(n)}(\bm{y})|$
$=$ $\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}\log\frac{\bar{f}_{1}^{(n)}(\bm{y})}{f_{1}^{(n)}(\bm{y})}\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty,$ (170)
and
$\underset{\bm{z}\in\mathbb{R}^{n}}{\sup}|\log\bar{h}_{1}^{(n)}(\bm{z})-\log h^{(n)}_{1}(\bm{z})|$
$=$ $\underset{\bm{z}\in\mathbb{R}^{n}}{\sup}\log\frac{\bar{h}_{1}^{(n)}(\bm{z})}{h^{(n)}_{1}(\bm{z})}\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty,$ (171)
which is equivalent to the following conclusions. For any $\bm{y},\bm{z}$ in $\mathbb{R}^{n}$
$\frac{\bar{f}_{1}^{(n)}(\bm{y})}{f_{1}^{(n)}(\bm{y})}\rightarrow 1\,\ \,\text{as}\,\,n\rightarrow\infty,$ (172)
$\frac{\bar{h}_{1}^{(n)}(\bm{z})}{h^{(n)}_{1}(\bm{z})}\rightarrow 1\,\ \,\text{as}\,\,n\rightarrow\infty.$ (173)
From (172) and (173), the conclusion of the lemma is obvious. ∎
Appendix D Proof of Lemma 7
Proof.
$\frac{W_{2}^{n}(\bm{z}|\bm{x})}{\bar{Q}^{(n)}_{1}(\bm{z})}$
$=$ $\frac{\frac{1}{(2\pi\sigma^{2})^{\frac{n}{2}}}e^{-\frac{\|\bm{z}-\bm{x}\|^{2}}{2\sigma^{2}}}}{\frac{1}{\big(2\pi(\sigma^{2}+\mu\Psi(n))\big)^{\frac{n}{2}}}e^{-\frac{\|\bm{z}\|^{2}}{2(\sigma^{2}+\mu\Psi(n))}}}$
$=$ $\bigg(\frac{\sigma^{2}+\mu\Psi(n)}{\sigma^{2}}\bigg)^{\frac{n}{2}}\cdot e^{\frac{-\mu\Psi(n)\|\bm{z}\|^{2}+2(\sigma^{2}+\mu\Psi(n))\bm{z}^{T}\bm{x}-(\sigma^{2}+\mu\Psi(n))\|\bm{x}\|^{2}}{2\sigma^{2}(\sigma^{2}+\mu\Psi(n))}}$
$=$ $\bigg(\frac{\sigma^{2}+\mu\Psi(n)}{\sigma^{2}}\bigg)^{\frac{n}{2}}e^{-\frac{\mu\Psi(n)\big\|\bm{z}-\frac{\sigma^{2}+\mu\Psi(n)}{\mu\Psi(n)}\bm{x}\big\|^{2}}{2\sigma^{2}(\sigma^{2}+\mu\Psi(n))}+\frac{\|\bm{x}\|^{2}}{2\mu\Psi(n)}}$
$\leq$ $\bigg(\frac{\sigma^{2}+\mu\Psi(n)}{\sigma^{2}}\bigg)^{\frac{n}{2}}e^{\frac{\|\bm{x}\|^{2}}{2\mu\Psi(n)}}$
$\leq$ $\bigg(\frac{\sigma^{2}+\mu\Psi(n)}{\sigma^{2}}\bigg)^{\frac{n}{2}}e^{\frac{n}{2\mu}}.$ (174)
Appendix E Proof of Lemma 10
Proof.
We have
[…]
$=$ $\bigg|\int Q^{(n)}_{\bm{Z},\bm{C}_{1}}\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{1}}}{Q^{(n)}_{\bm{Z},0}}d\bm{z}-\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{2}}}{Q^{(n)}_{\bm{Z},0}}d\bm{z}\bigg|$
$=$ $\bigg|\int Q^{(n)}_{\bm{Z},\bm{C}_{1}}\log Q^{(n)}_{\bm{Z},\bm{C}_{1}}d\bm{z}-\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log Q^{(n)}_{\bm{Z},\bm{C}_{2}}d\bm{z}$
$-\int(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}})\log Q^{(n)}_{\bm{Z},0}d\bm{z}\bigg|$
$=$ $\bigg|\int\big(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}}\big)\log Q^{(n)}_{\bm{Z},\bm{C}_{1}}d\bm{z}$
$+\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log Q^{(n)}_{\bm{Z},\bm{C}_{1}}d\bm{z}-\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log Q^{(n)}_{\bm{Z},\bm{C}_{2}}d\bm{z}$
$-\int(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}})\log Q^{(n)}_{\bm{Z},0}d\bm{z}\bigg|$
$\leq$ $\bigg|\int\big(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}}\big)\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{1}}}{Q^{(n)}_{\bm{Z},0}}d\bm{z}\bigg|$
$+$ $\bigg|\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{2}}}{Q^{(n)}_{\bm{Z},\bm{C}_{1}}}d\bm{z}\bigg|$ (175)
We analyze the two terms in (E) separately. For the first term, as the difference between $Q^{(n)}_{\bm{Z},\bm{C}_{1}}$ and $Q^{(n)}_{\bm{Z},\bm{C}_{2}}$ is the $k$-th component, we have
$\bigg|\int\big(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}}\big)\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{1}}}{Q^{(n)}_{\bm{Z},0}}d\bm{z}\bigg|$
$=$ $\frac{1}{M}\big|\int_{\bm{z}}\big(\frac{1}{(2\pi\sigma^{2})^{\frac{n}{2}}}e^{-\frac{\|\bm{z}-\bm{x}_{k}\|^{2}}{2\sigma^{2}}}-\frac{1}{(2\pi\sigma^{2})^{\frac{n}{2}}}e^{-\frac{\|\bm{z}-\bm{x}_{k}^{\prime}\|^{2}}{2\sigma^{2}}}\big)$
$\times$ $\log\frac{1}{M}\sum_{i=1}^{M}e^{\frac{\|\bm{z}\|^{2}-\|\bm{z}-\bm{x}_{i}\|^{2}}{2\sigma^{2}}}d\bm{z}\big|$
$\overset{(a)}{\leq}$ $\frac{1}{M}\int_{\bm{z}}\big(\frac{1}{(2\pi\sigma^{2})^{\frac{n}{2}}}e^{-\frac{\|\bm{z}-\bm{x}_{k}\|^{2}}{2\sigma^{2}}}+\frac{1}{(2\pi\sigma^{2})^{\frac{n}{2}}}e^{-\frac{\|\bm{z}-\bm{x}_{k}^{\prime}\|^{2}}{2\sigma^{2}}}\big)$
$\times$ $\log e^{\frac{n\Psi(n)}{2\sigma^{2}}}d\bm{z}$
$=$ $\frac{2}{M}\frac{n\Psi(n)}{2\sigma^{2}}\log e$
$=$ $\frac{n\Psi(n)}{M\sigma^{2}}\log e$ (176)
where step (a) follows by using
$\|\bm{x}\|^{2}\leq\|\bm{z}\|^{2}-\|\bm{z}-\bm{x}_{i}\|^{2}\leq\|\bm{x}\|^{2}$ (177)
and each $\bm{x}$ satisfies $\mu n\Psi(n)\leq\|\bm{x}\|^{2}\leq n\Psi(n)$.
For the second term, let
$p_{1}=W_{2}^{n}(\bm{Z}|\bm{x}_{k}),\,\ \ \ \,q_{1}=W_{2}^{n}(\bm{Z}|\bm{x}_{k}^{\prime})$ (178)
and
$p_{2}=q_{2}=\frac{1}{M-1}\sum_{i=1,i\neq k}^{M}W_{2}^{n}(\bm{Z}|\bm{x}_{i}),$ (179)
then $Q^{(n)}_{\bm{Z},\bm{C}_{1}}$ and $Q^{(n)}_{\bm{Z},\bm{C}_{2}}$ can be expressed as
$Q^{(n)}_{\bm{Z},\bm{C}_{1}}=\frac{1}{M}p_{1}+\frac{M-1}{M}p_{2},$
$Q^{(n)}_{\bm{Z},\bm{C}_{2}}=\frac{1}{M}q_{1}+\frac{M-1}{M}q_{2}.$ (180)
We have
$\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{2}}}{Q^{(n)}_{\bm{Z},\bm{C}_{1}}}d\bm{z}$
$=$ $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{2}}\|Q^{(n)}_{\bm{Z},\bm{C}_{1}})$
$=$ $\mathbb{D}(\frac{1}{M}q_{1}+\frac{M-1}{M}q_{2}\|\frac{1}{M}p_{1}+\frac{M-1}{M}p_{2})$
$\overset{(b)}{\leq}$ $\frac{1}{M}\mathbb{D}(q_{1}\|p_{1})+\frac{M-1}{M}\mathbb{D}(q_{2}\|p_{2})$
$=$ $\frac{1}{M}\mathbb{D}(q_{1}\|p_{1})$
$\overset{(c)}{=}$ $\frac{\|\bm{x}_{k}-\bm{x}^{\prime}_{k}\|^{2}}{2M\sigma^{2}}$
$\overset{(d)}{\leq}$ $\frac{2n\Psi(n)}{M\sigma^{2}}$ (181)
where (b) holds because KL divergence $\mathbb{D}(p\|q)$ is a convex function with the pair $(p,q)$ [38], (c) follows from the close form of KL divergence between two Gaussian distributions [39], and (d) follows from the fact both $\|\bm{x}_{k}\|^{2}$ and $\|\bm{x}^{\prime}_{k}\|^{2}$ are in $\mu^{2}n\Psi(n),n\Psi(n)]$. Combining (E), (E) and (E), we conclude
$\bigg|\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{1}}\|Q^{(n)}_{\bm{Z},0})-\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{2}}\|Q^{(n)}_{\bm{Z},0})\bigg|\leq\frac{4n\Psi(n)}{M\sigma^{2}}.$ (182)
Appendix F Proof of Theorem 9
The TVD between $P_{1}$ and $P_{0}$ is written as
$\begin{split}&\mathbb{V}(P_{1},P_{0})\\ =&\frac{1}{2}\int_{\mathbb{R}^{n}}|p_{1}(\bm{x})-p_{0}(\bm{x})|d\bm{x}\\ =&\frac{\varepsilon}{2}\int_{\mathbb{R}^{n}}p_{0}(\bm{x})|h_{\varepsilon}(\bm{x})|d\bm{x}\\ \leq&\frac{\varepsilon}{2}\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h_{\varepsilon}(\bm{x})|\\ =&\,\ \,\frac{1}{2}\varepsilon.\end{split}$ (183)
Then KL divergence between $P_{1}$ and $P_{0}$ is written as
$\mathbb{D}(P_{1}\|P_{0})$
$=$ $\int_{\mathbb{R}^{n}}p_{1}(\bm{x})\log\frac{p_{1}(\bm{x})}{p_{0}(\bm{x})}d\bm{x}$
$=$ $\int_{\mathbb{R}^{n}}p_{1}(\bm{x})\log\left(1+\varepsilon h_{\varepsilon}(\bm{x})\right)d\bm{x}$
$\overset{(a)}{=}$ $\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\left(1+\varepsilon h_{\varepsilon}(\bm{x})\right)$
$\times$ $\left[\epsilon h_{\varepsilon}(\bm{x})-\frac{1}{2}\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})\right]d\bm{x}+o(\varepsilon^{2})$
$=$ $\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\left[\varepsilon h_{\varepsilon}(\bm{x})+\frac{1}{2}\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})\right$.
- $\left.\frac{1}{2}\varepsilon^{3}h_{\varepsilon}^{3}(\bm{x})\right]d\bm{x}+o(\varepsilon^{2})$
$\overset{(b)}{=}$ $\frac{1}{2}\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})d\bm{x}+o(\varepsilon^{2})$
$\overset{(c)}{\leq}$ $\,\,\frac{1}{2}\varepsilon^{2}\log e+o(\varepsilon^{2})$
$\overset{(d)}{\leq}$ $\,\,\varepsilon^{2}\log e$ (184)
In the above inequality, we have used Taylor’s theorem in (a), Lemma 14, $\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h_{\varepsilon}(\bm{x})|\leq 1$ and the integrability assumption in in (b) and (c) and $\varepsilon$ is a small number in (d).
Appendix G Proof of Theorem 10
Let $\bm{x}=(x_{1},\cdots,x_{n})$ be a random vector and each coordinate is independent. As $\|\cdot\|_{w}^{*}$ is a metric, with a slight abuse of notion, we have
[…]
We consider the two terms in the equation (185) separately. For the first part, using the characteristic function of Gaussian distribution and Lemma 15, the characteristic functions of $P_{n}^{(n)}(x_{1},\cdots,x_{n})$ and $P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})\cdot\delta_{0}^{1}(x_{n})$ are expressed as
$\exp\left[-\frac{1}{2}\mu\Psi(n)\cdot(t_{1}^{2}+\cdots+t_{n}^{2})\right]$
and
$\exp\left[-\frac{1}{2}\mu\Psi(n)\cdot\cdot(t_{1}^{2}+\cdots+t_{n-1}^{2})\right].$
Hence, we have
$\Big|\exp\left[-\frac{1}{2}\mu\Psi(n)\cdot(t_{1}^{2}+\cdots+t_{n}^{2})\right]$
- $\exp\left[-\frac{1}{2}\mu\Psi(n)\cdot(t_{1}^{2}+\cdots+t_{n-1}^{2})\right]\Big|$
$\leq$ $\exp\left[-\frac{1}{2}\mu\Psi(n)\cdot(t_{1}^{2}+\cdots+t_{n-1}^{2})\right]$
$\times$ $\big|\exp\left(-\frac{1}{2}\mu\Psi(n)t_{n}^{2}\right)-1\big|$
$\leq$ $\big|\exp\left(-\frac{1}{2}\mu\Psi(n)t_{n}^{2}\right)-1\big|.$ (186)
Since $t_{n}$ is a scale in $\mathbb{R}$ and is independent with $\Psi(n)$, we have $\frac{1}{2}\mu\Psi(n)\cdot t_{n}^{2}=O(\frac{\mu\cdot t_{n}^{2}}{n^{\tau}})\rightarrow 0$ as $n\rightarrow\infty$, which means (G) tends to 0 uniformly. Hence, we have $\|P_{n}^{(n)}(x_{1},\cdots,x_{n})-P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})\cdot\delta_{0}^{1}(x_{n})\|_{w}^{*}\rightarrow 0$.
For the second part, as the Dirac measure $\delta_{0}^{n}(x_{1},\cdots,x_{n})$ can be written as $\prod_{i=1}^{n}\delta_{0}^{1}(x_{i})$, and the random variable $x_{n}$ in both terms $P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})\cdot\delta_{0}^{1}(x_{n})$ and $\delta_{0}^{n}(x_{1},\cdots,x_{n})$ are the same Dirac measure and are independent with the other random variable $x_{1}$ to $x_{n-1}$, we have
$\|P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})\cdot\delta_{0}^{1}(x_{n})-\delta_{0}^{n}(x_{1},\cdots,x_{n})\|_{w}^{*}$ (187)
is equivalent to
$\|P_{n-1}^{(n)}(x_{1},\cdots,x_{n-1})-\delta_{0}^{n-1}(x_{1},\cdots,x_{n-1})\|_{w}^{*}.$ (188)
Using the same skill as (G), we have
[…]
Note that two terms in the equation in (189) are n1n-1 dimension situation as (185). Using the technique repeatedly, finally we just need to prove $\|P_{1}^{(n)}(x_{1})-\delta_{0}^{1}(x_{1})\|_{w}^{*}\rightarrow 0$ as $n\rightarrow\infty$. The characteristic functions of $P_{1}^{(n)}(x_{1})$ and $\delta_{0}^{1}(x_{1})$ are $\exp\left(-\frac{1}{2}\mu\Psi(n)t_{1}^{2}\right)$ and $1$, respectively. Following the same arguments above, we have $\|P_{1}^{(n)}(x_{1})-\delta_{0}^{1}(x_{1})\|_{w}^{*}\rightarrow 0$.
Appendix H Proof of Lemma 17
Let $\sigma_{1}^{2}=1+\theta_{n}$, we have the square of Hellinger distance of two Gaussian distribution can be expressed as [39]
$H^{2}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})=1-\left(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}}\right)^{\frac{n}{2}}.$ (190)
To prove (158), we just need to prove
$(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}\rightarrow\begin{cases}1&\tau>\frac{1}{2},\\ 0&0<\tau<\frac{1}{2},\end{cases}\,\ \,n\rightarrow\infty,$ (191)
which is equivalent to
$\frac{n}{2}\cdot\ln(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})\rightarrow\begin{cases}0&\tau>\frac{1}{2},\\ -\infty&0<\tau<\frac{1}{2},\end{cases}\,\ \,n\rightarrow\infty.$ (192)
The left-hand of (192) can be formulated as follows,
$\frac{1}{2}n\ln\frac{2\sigma_{1}}{1+\sigma_{1}^{2}}$
$=$ $\frac{1}{2}n\ln\frac{2(1+\theta_{n})^{\frac{1}{2}}}{(2+\theta_{n})}$
$=$ $\frac{1}{4}n\ln\bigg(\frac{2(1+\theta_{n})^{\frac{1}{2}}}{(2+\theta_{n})}\bigg)^{2}$
$=$ $\frac{1}{4}n\ln\frac{4(1+\theta_{n})}{(2+\theta_{n})^{2}}$
$=$ $\frac{1}{4}n\ln\frac{4+4cn^{-\tau}}{c^{2}n^{-2\tau}+4cn^{-\tau}+4}$
$=$ $\frac{1}{4}n\ln\left[1-\frac{c^{2}n^{-2\tau}}{c^{2}n^{-2\tau}+4cn^{-\tau}+4}\right]$
$\sim$ $-\frac{1}{4}\cdot\frac{c^{2}n^{1-2\tau}}{c^{2}n^{-2\tau}+4cn^{-\tau}+4}$ (193)
  • When $\tau>\frac{1}{2}$, it is obvious that the term (H) approaches 0 from the negative axis. Thus, we have proved the first part of (192).
  • When $\tau<\frac{1}{2}$, $1-2\tau>0$ and the term (H) approaches -$\infty$ as $n\rightarrow\infty$. Thus, the second part of (192) is proved.